### v1

# Galois Action and Localization in Number Fields

Abstract.
For a Galois number field $K$, the Galois group $\text{Gal}(K/\mathbb{Q})$ acts on the class group $Cl_{K}$ in a very natural way: $\sigma\cdot[I]=[\sigma(I)]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$, $I]\in Cl_{K}$. In this paper, we will explore how the unique properties of this group action work together to elucidate the relationship between these two groups. While previous work on this problem has focused on representation theory, we take a direct approach to some classical and new problems. The paper concludes with an exploration of the class groups of localizations of the ring of integers $\mathcal{O}_{K}$. These turn out to be powerful tools for understanding $Cl_{K}$ and overrings of $\mathcal{O}_{K}$.
1. Introduction and Notation
Throughout this paper $K$ will denote a Galois number field with Galois group $G:=\text{Gal}(K/\mathbb{Q})$, and $\mathcal{O}_{K}$ its ring of integers with class group $Cl_{K}$ and class number $h_{K}=|Cl_{K}|$. The Galois group of $K$ acts on the class group in a very natural way. For any $\sigma\in G$ and $I]\in Cl_{K}$, we can define the action $\sigma\cdot[I]=[\sigma(I)]$. This is a well-defined group action which provides us with some tools for characterization the relationship between the two groups. More than this, we also have the peculiar property that $\sigma\cdot([I][J])=(\sigma\cdot[I])(\sigma\cdot[J])$ for any $\sigma\in G$ and $I],[J]\in Cl_{K}$. This induces a map from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\mapsto\bar{\sigma}$ where $\bar{\sigma}([I])=[\sigma(I)]$. Finally, we have the “norm property” that $\prod_{\sigma\in G}\sigma\cdot[I]=\left[\prod_{\sigma\in G}\sigma(I)\right]=\text{Prin}(\mathcal{O}_{K})$ for any $I]\in G$—for details, see [12]. Note, in this paper we will study the case where $K$ is Galois over $\mathbb{Q}$ with the knowledge that the theorems herein apply also to the case when $K/L$ is a Galois extension of number fields with $\mathcal{O}_{L}$ a principal ideal domain (PID). Now, these characteristics of the action motivate the following definition.
Definition 1.1.
Let $G$ and $A$ be groups with $A$ abelian, and let
$\alpha:G\times A\longrightarrow A,\quad(g,a)\mapsto g\cdot a$
be a map. If $\alpha$ satisfies the following properties:
  1. (1)
    $g_{1}\cdot(g_{2}\cdot a)=(g_{1}g_{2})\cdot a$ for all $g_{1},g_{2}\in G$, $a\in A$,
  2. (2)
    $e_{G}\cdot a=a$ for all $a\in A$,
  3. (3)
    $g\cdot(a_{1}a_{2})=(g\cdot a_{1})(g\cdot a_{2})$ for all $g\in G$, $a_{1},a_{2}\in A$,
  4. (4)
    $\prod_{g\in G}(g\cdot a)=e_{A}$ for all $a\in A$,
then we say that $\alpha$ is a norm-like action.
Throughout this paper, we will see how properties 1-4 can be used to place strong restrictions on the structure of $Cl_{K}$ given $G$ and vice versa. We will place a particular emphasis on techniques developed and some applications to factorization theory.
The previous work done on this subject has been primarily through the lens of representation theory—considering $Cl_{K}$ as a $G$-module. This approach was utilized by Fröhlich in 1952 ([8]) and subsequently many others. Cornell and Rosen used this approach in [5] to give results on the structure of $Cl_{K}$ when $h_{K}$ is known. Others such as Lemmermeyer ([11]) and Iwasawa ([10]) used the $G$-module structure of $Cl_{K}$ along with assumptions on intermediate fields to give various results on the $p$-rank of the class group. For a thorough review of the previous work done in this area, see [13].
In the following section, we will take a closer look at the conditions of Definition 1.1 and their implications for norm-like actions in general. In Section 3, we explore how conditions 1-2, together with 4, place restrictions of the structure of $Cl_{K}$ for a Galois number field of given degree. In Section 4, we focus on applications of condition 3 to the inverse class group problem. In particular, which number fields $K$ can have a given abelian group as their class group. Finally, in Section 5, the authors determine the structure of the class group for localizations of the form $\mathcal{O}_{K}[\frac{1}{x}]$ which ultimately strengthens many of the previously developed results and leads to some interesting observations about rings of integers.
2. Norm-Like Group Actions
In this section, we will consider norm-like group actions directly. The results and observation made here will be used throughout the remainder of the paper. First, let us recall the definitions of a norm-like action.
Definition 1.1.
Let $G$ and $A$ be groups with $A$ abelian, and let
$\alpha:G\times A\longrightarrow A,\quad(g,a)\mapsto g\cdot a$
be a map. If $\alpha$ satisfies the following properties:
  1. (1)
    $g_{1}\cdot(g_{2}\cdot a)=(g_{1}g_{2})\cdot a$ for all $g_{1},g_{2}\in G$, $a\in A$,
  2. (2)
    $e_{G}\cdot a=a$ for all $a\in A$,
  3. (3)
    $g\cdot(a_{1}a_{2})=(g\cdot a_{1})(g\cdot a_{2})$ for all $g\in G$, $a_{1},a_{2}\in A$,
  4. (4)
    $\prod_{g\in G}(g\cdot a)=e_{A}$ for all $a\in A$,
then we say that $\alpha$ is a norm-like action.
Conditions $1$ and $2$ are those of a typical group action. The definition diverges with condition $3$ which ensures the action of each element $g$ on $A$ behaves like a homomorphism. In fact, each element of $G$ can be identified with an endomorphism of $A$ via the map $g\mapsto\phi$ where $\phi(a)=g\cdot a$. If $A$ is finite, in the case where $g\cdot a=e_{A}\implies a=e_{A}$, this map defines a group homomorphism from $G$ to $\text{Aut}(A)$, the kernel of which is the set $\{g\in G\,|\,g\cdot a=a\,\forall\,a\in A\}$. The implications of this are closely considered in Section 4. We also note that condition $3$ guarantees $g\cdot e_{A}=e_{A}$ for all $g\in G$. Thus, a norm-like group action will be transitive if and only if $A$ is trivial.
Let us consider a group action of $G$ on $A$ satisfying conditions 131-3. First, it suffices to check condition $4$ for the generators of $A$. Now, consider the subgroup $H:=\{\prod_{g\in G}g\cdot a\,|\,a\in A\}\unlhd A$. For any $x\in H$, we have $x=\prod_{g\in G}g\cdot a^{\prime}$ for some $a^{\prime}\in A$. Thus, for any $g^{\prime}\in G$, $g^{\prime}\cdot x=g^{\prime}\cdot\left(\prod_{g\in G}g\cdot a^{\prime}\right)=\prod_{g\in G}(g^{\prime}*g)\cdot a^{\prime}=\prod_{g\in G}g\cdot a^{\prime}=x$. Hence, $H$ is a subgroup of the elements of $A$ with trivial orbit. Therefore, if $e_{A}$ is the only element with a trivial orbit, condition $4$ follows from 131-3.
Now, let $A$ be an arbitrary abelian group, and $G=\{id,-id\}\leqslant\text{Aut}(A)$ where $id$ is the identity, and id(a)=a-id(a)=-a for all $a\in A$, which is a automorphism just in case $A$ is abelian. It is easy to verify that the natural action $g\cdot a=g(a)$ is norm-like. This case has been of particular interest over the last three years as the study of weighted zero-sum sequences has become popular in connection with normset factorization (see [1] and [7]). In particular, $\{id,-id\}$ is the subgroup induced by the action of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ when $K$ is quadratic. As we have seen, $|G|=2$ places no restriction on $A$. As we will see throughout this paper, this is a very unique case. Even $|G|=3$ is much more restrictive and the relevant arithmetic far more complicated.
3. Direct Consequences of the Group Action
In this section, we produce results on the structure of the class group of Galois number fields which follow directly from the action of the Galois group $G=\text{Gal}(F/\mathbb{Q})$ on the class group $Cl_{K}$ in conjunction with the norm property $\prod_{\sigma\in G}\sigma(I)\in\text{Prin}(\mathcal{O}_{K})$ for any ideal $I\subseteq\mathcal{O}_{K}$. We begin with a lemma which we will use extensively throughout this paper.
Lemma 3.1.
Let $K$ be a Galois number field with $G=\text{Gal}(K/\mathbb{Q})$ and class group $Cl_{K}$. If $G$ acts trivially on $I]\in Cl_{K}$, the order of $[I]$ divides $|G|=[K:\mathbb{Q}]$.
Proof.
Assume $G$ acts trivially on $I]\in Cl_{K}$. Now, by the norm property, we have
$\prod_{\sigma\in G}\sigma(I)\in\text{Prin}(\mathcal{O}_{K}).$
Thus, we have
$\prod_{\sigma\in G}[\sigma(I)]=\text{Prin}(\mathcal{O}_{K})\Rightarrow{}\prod_{\sigma\in G}\sigma\cdot[I]=\text{Prin}(\mathcal{O}_{K}).$
As $G$ acts trivially on $[I]$, we have
$\text{Prin}(\mathcal{O}_{K})=\prod_{\sigma\in G}\sigma\cdot[I]=\prod_{\sigma\in G}[I]=[I]^{|G|}.$
This completes the proof. ∎
The first result we present follows from Lemma 2 in [5] using the representation theoretic approach. We will prove it directly here and generalize the result in section $4$ with the techniques developed therein.
Theorem 3.2.
Let $K$ be a Galois number field of degree $p^{r}$. Then, $h_{K}\equiv 0\text{ or }1\text{ mod }p$.
Proof.
Assume that $h_{K}\not\equiv 1$ mod $p$. Once again, let $G$ act on $Cl_{K}$ via the action $\sigma\cdot[I]=[\sigma(I)]$. By the orbit-stabilizer theorem, the lengths of the orbits of this group action divide $|G|=[K:\mathbb{Q}]=p^{r}$. In particular, they must be elements of the set $\{1,p,p^{2},\dots,p^{r}\}$.
Now, the identity of $Cl_{K}$ has a trivial orbit of order $1$. The orbits partition $Cl_{K}$, and all non-trivial orbits have order congruent to 0 mod $p$. Hence, as we assumed $h_{K}\not\equiv 1$ mod $p$, there must be some non-identity class of $Cl_{K}$ in a trivial orbit. Call this element $J]\neq\text{Prin}(\mathcal{O}_{K})$. By Lemma 3.1, $[J]$ has order dividing $p^{r}$, and as we assumed $[J]$ was not the identity, $p$ must divide the order of $[J]$. Therefore, by Lagrange’s theorem, $p\text{ divides }h_{K}$, so $h_{K}\equiv 0$ mod $p$. This completes the proof.
This theorem demonstrates one of the many ways in which quadratic number fields are exceptional. Beyond their peculiarity of being Galois in general, Theorem 3.2 leaves all class numbers permissible in the quadratic case. The same is true of the following.
Theorem 3.3.
If $K$ is a Galois number field of degree $n$ and $p$ the smallest prime divisor of $n$, $h_{k}=1$ or $h_{k}\geqslant p$.
Proof.
Assume for the purpose of contradiction that $1<h_{k}<p$. As $h_{k}<p$, the smallest prime divisor of $|G|$, the orbit-stabilizer theorem tells us every element of $Cl_{K}$ has a trivial orbit under the action of $G$ on $Cl_{K}$. As $h_{k}>1$, we may take $J]\neq\text{Prin}(\mathcal{O}_{K})$ in $Cl_{K}$. Then, by Lemma 3.1, we must have $J]^{|G|}=\text{Prin}(\mathcal{O}_{K})$, so the order of $[J]$ divides $|G|=n$. However, this is a contradiction as $1<|[J]|<h_{k}<p$, and $p$ is the smallest prime divisor of $n$. ∎
We now turn our attention to some results on Galois number fields of odd degree.
Theorem 3.4.
Let $K$ be a Galois number field of odd degree. Then, $Cl_{K}$ cannot have a unique involution.
Proof.
Assume for the purpose of contradiction that $Cl_{K}$ has a unique element of order 2, call it $[J]$. Thus, as each $\sigma\in G$ is an automorphism, we must have $\sigma\cdot[J]=[\sigma(J)]=[J]$ for all $\sigma\in G$. Hence, $G$ acts trivially on $[J]$, so by Lemma 3.1 the order of $[J]$ divides $|G|$. However, this is a contradiction as $[J]$ has order $2$ and $|G|$ is odd. ∎
In particular, for a Galois number field of odd degree, $Cl_{K}$ cannot have a unique invariant factor of even order. Notably, this precludes cyclic groups of even order. Now, from Theorem 3.4 we get the following factorization result which first appears as Lemma 3.5 in [6].
Corollary 3.5.
Let $K$ be a Galois number field of odd degree. Then $\mathcal{O}_{K}$ is an HFD if and only if $\mathcal{O}_{K}$ is a UFD.
Proof.
A well-known result from Carlitz ([2]) tells us $\mathcal{O}_{K}$ is an HFD if and only if $h_{K}=1$ or $2$. Noting that Theorem 3.4 disallows $h_{K}=2$ and $\mathcal{O}_{K}$ is a UFD if and only if $h_{K}=1$ completes the proof. ∎
Note that this result also follows directly from Theorem 3.3. We conclude this section with a specific application of these results.
Theorem 3.6.
Let $p<23$ be an odd prime and $a\in\mathbb{Z}$ not divisible by the $p^{th}$ power of any prime. If $K$ is the splitting field of $x^{p}-a$, $\mathcal{O}_{K}$ is an HFD if and only if it is $\mathcal{O}_{K}$ is a UFD.
Proof.
We first note that, as the splitting field of $x^{p}-a$, $K$ is Galois over $\mathbb{Q}$. Furthermore, we have $K\cong\mathbb{Q}(\omega,\alpha)$ where $\omega$ is a primitive $p^{th}$ root of unity and $\alpha$ a root of $x^{p}-a$. This gives us the following lattice.
$\mathbb{Q}$$\mathbb{Q}(\omega)$$\mathbb{Q}(\alpha)$$\mathbb{Q}(\omega,\alpha)$
Where $\mathbb{Q}(\omega,\alpha)$ is Galois over $\mathbb{Q}(\omega)$ of degree $p$. Now, as $p<23$, $\mathbb{Z}[\omega]$ is a UFD, so the result follows from Corollary 3.5.
[…]
Recall that the action of $G=\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ induces a homomorphism from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\overset{\psi}{\mapsto}\phi$ where $\phi([I])=\sigma\cdot[I]=[\sigma(I)]$. In this section, we will see how this map can be used to further restrict the structure of $Cl_{K}$ for a Galois number field $K$. Conversely, we will use this tool to help answer questions related to the inverse class group problem. This terminology is used in the spirit of the inverse Galois problem. In particular, the inverse class group problem asks if an arbitrary abelian group can be realized as the class group of a ring of integers. This question was famously answered in the positive for the more general class of Dedekind domains by Claborn in [4]. In the case of imaginary quadratics, the answer is known to be no. Certain groups can be shown by brute force not to appear, the smallest among them being $(\mathbb{Z}/3\mathbb{Z})^{3}$. For details, the reader is encouraged to see (CITE). The question remains open for rings of integers in general. In this section, for a given finite abelian group $A$, we will explore which rings of integers may admit $Cl_{K}\cong A$. We begin with a direct proof of another result from [5] which we will return to with some new found tools in section 4.
Theorem 4.1.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, $Cl_{K}\ncong\mathbb{Z}/p^{n}\mathbb{Z}$ for $n\geqslant 2$.
Proof.
[…]
Assume $H$ is trivial. Then, each element of $G$ acts trivially on the class group. Therefore, by Lemma 3.1, every element of $Cl_{K}$ has order dividing $|G|=p$, but this contradicts our assumption that $Cl_{K}\cong\mathbb{Z}/p^{n}\mathbb{Z}$ with $n\geqslant 2$. Hence, $H$ is the unique subgroup of order $p$.
[…]
$\prod_{\sigma\in G}\sigma(J)\in\text{Prin}(\mathcal{O}_{K})$
implies
$\prod_{\sigma\in G}[\sigma(J)]=\prod_{\sigma\in G}\sigma[(J)]=\prod_{i=0}^{p-1}\psi^{i}([J])=\text{Prin}(\mathcal{O}_{K})$
Furthermore, $\psi^{i}([J])=(\gamma^{(p-1)p^{n-2}})^{i}([J])=[J]^{\left(\alpha^{(p-1)p^{n-2}}\right)^{i}}$. Note, $(p-1)p^{n-2}=\varphi(p^{n-1})$, so $\alpha^{(p-1)p^{n-2}}\equiv 1$ mod $p^{n-1}$. Also, because $\alpha$ is a primitive root mod $p^{n}$, each $(\alpha^{(p-1)p^{n-2}})^{i}=\alpha^{i(p-1)p^{n-2}}$ is unique mod $p^{n}$ for each $0\leqslant i\leqslant p-1$. Hence, they are precisely the elements $\{1,p^{n-1}+1,2p^{n-1}+1,\dots,(p-1)p^{n-1}+1\}$ mod $p^{n}$. Taking their sum, we get
$\sum_{i=0}^{p-1}\left(1+i\cdot p^{n-1}\right)=p+p^{n-1}\sum_{i=1}^{p-1}i=p+p^{n-1}\cdot\frac{(p-1)p}{2}=p+\frac{p-1}{2}\cdot p^{n}\equiv p\text{ mod }p^{n}$
Hence,
$\prod_{i=0}^{p-1}\psi^{i}([J])=\prod_{i=0}^{p-1}[J]^{\left(\alpha^{(p-1)p^{n-2}}\right)^{i}}=[J]^{\sum_{i=0}^{p-1}1+i\cdot p^{n-1}}=[J]^{p}=\text{Prin}(\mathcal{O}_{K})$
But this implies the order of $[J]$ divides $p$, and $[J]$ is of order $p^{n}$ with $n\geqslant 2$. This is a contradiction. Therefore, $Cl_{K}$ cannot be $\mathbb{Z}/p^{n}\mathbb{Z}$ for any $n\geqslant 2$.
Recall that Theorem 3.2 tells us (in particular) that for a Galois number field of degree $p$, we must have $h_{k}\equiv 0$ or $1$ mod $p$. Thus, Theorem 4.1 gives us a further restriction in the first case on the structure of the class group. We now shift our attention to the inverse class group. First, we consider the case when $h_{k}$ is assumed to be an odd prime rather than $K:\mathbb{Q}]$. Note, this is equivalent to asking which number fields $K$ can have $Cl_{K}\cong\mathbb{Z}/p\mathbb{Z}$.
Theorem 4.2.
Let $K$ a Galois number field with $n=[K:\mathbb{Q}]$ and $h_{K}=p$ prime. Then, $p|n$ or gcd(p1,n)>1(p-1,n)>1.
Proof.
[…]
This theorem tells us that having class group $\mathbb{Z}/p\mathbb{Z}$ is a relatively restrictive condition. For example, only Galois number fields of degree divisible by $2$ or $17$ can have class group $\mathbb{Z}/17\mathbb{Z}$. We continue this section with a few specific examples of the inverse class group problem which present new techniques and highlight the interesting interplay between the induced homomorphism and the norm property of the group action.
Example 4.3.
Let us assume $Cl_{K}\cong\mathbb{Z}/13\mathbb{Z}$ for some Galois number field $K$. None of the theorems developed thus far preclude a cubic Galois extension from having such a class group, and indeed we see that the number field with defining polynomial $x^{3}-x^{2}-354x-2441$ is one such example. Now, any cubic Galois extension has $G\cong\mathbb{Z}/3\mathbb{Z}$, so the induced homomorphism $\psi:G\rightarrow\text{Aut}(Cl_{K})$ must be injective as the alternative would imply there exists an element in $\mathbb{Z}/13\mathbb{Z}$ of order $|G|=3$.
[…]
The following theorem is inspired by this example and demonstrates that the phenomenon described at the end must occur in general for $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$.
Theorem 4.4.
Let $p$ be prime and $\phi=(a_{1},a_{2},...,a_{p-1})$ a cyclic generator of $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$. If $n$ properly divides p1p-1, then the elements in each disjoint cycle of $\phi^{n}$ sum to 0 mod $p$.
Proof.
Say $\phi(1)=a$, then we must have $\phi(x)=a\cdot x$ for any $x\in\mathbb{Z}/p\mathbb{Z}$. This also implies that $a$ is a primitive root mod $p$. Now, any $x\in\mathbb{Z}/p\mathbb{Z}$ is contained in a (p1)/n(p-1)/n cycle in $\phi^{n}$. Thus, the cycle is of the form $(x\,\,\,xa^{n}\,\,\,xa^{2n}\dotsm xa^{(\frac{p-1}{n}-1)n})$. Taking the sum, we get
$x\sum_{r=0}^{\frac{p-1}{n}-1}(a^{n})^{r}=x\cdot\frac{1-(a^{n})^{\frac{p-1}{n}}}{1-a^{n}}=x\cdot\frac{1-a^{p-1}}{1-a^{n}}$
By Fermat’s Little Theorem, we have $1-a^{p-1}\equiv 0$ mod $p$. The result then follows from the fact that $n$ properly divides p1p-1, so $1-a^{n}\not\equiv 0$ mod $p$.
Example 4.5.
Consider a Galois number field $K$ with class group $Cl_{K}\cong(\mathbb{Z}/2\mathbb{Z})^{3}$. It is well known that $\text{Aut}((\mathbb{Z}/2\mathbb{Z})^{3})\cong\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$, the group of invertible $3\times 3$ matrices over the field of two elements with order $|\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})|=168=2^{3}\cdot 3\cdot 7$. Employing the same methods as in the proof of Theorem 4.2, we immediately see that the order of the extension $n=[K:\mathbb{Q}]$ must be divisible by $2,3$ or $7$. However, Theorem 3.2 tells us that no Galois extension of degree $3^{r}$ can have class group $(\mathbb{Z}/2\mathbb{Z})^{3}$ for any $r\geqslant 1$. We will now show that in fact no Galois number field of degree $3m$ where $\text{gcd}(2,m)=\text{gcd}(7,m)=1$ can have class group $(\mathbb{Z}/2\mathbb{Z})^{3}$.
Assume $G=3m$ where $m$ is not divisible by $2$ or $7$. Once again, the homomorphism $\psi:G\rightarrow\text{Aut}(Cl_{K})\cong\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ must be non-trivial as $Cl_{K}$ has no non-identity elements of order dividing $|G|=3m$. Thus, $im(\psi)$ must be a subgroup of order $3$ by the first isomorphism theorem. Now, consider the following element of order $3$ in $\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ and its powers:
$E=\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&1&1\end{bmatrix}\hskip 14.22636ptE^{2}=\begin{bmatrix}1&0&0\\ 0&1&1\\ 0&1&0\end{bmatrix}\hskip 14.22636ptE^{3}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
Now, $E$ has characteristic polynomial $p_{E}(x)=(x+1)(x^{2}+x+1)$, so we see it has eigenvalue $\lambda=1$. From here, one can show $(1,0,0)^{T}$ is a corresponding eigenvector. That is, $E$ fixes the vector $(1,0,0)^{T}$, and so $E^{2}$ must also. Furthermore, by the second Sylow theorem and the fact that similar matrices have the same eigenspaces, we conclude that any element of order $3$ in $\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ fixes $(1,0,0)^{T}$. Therefore, as $im(\psi)$ is a subgroup of order $3$, we must have $(1,0,0)\in(\mathbb{Z}/2\mathbb{Z})^{3}$ fixed by every element of $G$, but this implies the order of $(1,0,0)$, namely $2$, divides $|G|=3m$, a contradiction. Therefore, we see that some degrees such as $K:\mathbb{Q}]=15$ which were not previously disallowed are impossible for $Cl_{K}\cong(\mathbb{Z}/2\mathbb{Z})^{3}$.
Similar methods can be employed to investigate the inverse class group problem for $p$-elementary abelian groups as well as cyclic groups of order $2^{m}p^{r}$ with $m\in\{0,1\}$ and $r\geqslant 1$ where $p$ is an odd prime.
5. Localizations of Rings of Integers
In this section, we will use localization to strengthen some of the results developed previously. First, recall that any overring—in particular, any localization—of a Dedekind domain remains Dedekind (for further information see CITE). Intuitively, as a localization of $R$ is attained by turning a set of elements $S\subseteq R$ into units, we expect the class group of $R_{S}$ to be smaller than that of $R$ as those non-principal ideals $I\subseteq R$ for which $I\cap S\neq\emptyset$ will become principal. As we will demonstrate, “small” localizations will tend to lead to small changes in the class group. In the context of rings of algebraic integers, this will be useful in further restricting the possible structure of $Cl_{K}$.
Theorem 5.1.
Let $D$ be a Dedekind domain and $x\in D$ a nonzero nonunit with
$(x)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}$
Then $Cl(D[\frac{1}{x}])\cong Cl(D)/\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>.
Note, $D[\frac{1}{x}]$ is the localization of $D$ at the multiplicative set $\{x^{n},\,n\in\mathbb{N}\}$.
Proof.
Let $\phi:Cl(D)\rightarrow Cl(D[\frac{1}{x}])$ be the canonical map $I]\overset{\phi}{\mapsto}\left[ID[\frac{1}{x}]\right]$. First we will show that $\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> is contained in $\text{ker}(\phi)$. It suffices to show that $\mathfrak{p}_{i}$ becomes principal in $D[\frac{1}{x}]$ for any $1\leqslant i\leqslant k$. This is clearly the case as $x\in\mathfrak{p}_{i}$, so $\mathfrak{p}_{i}D[\frac{1}{x}]=D[\frac{1}{x}]$.
Now, let $I]\in\text{ker}(\phi)$. The case when $I]=\text{Prin}(D)$ is trivial, so we will assume $I$ is a non-principal, integral ideal. As $D$ is Dedekind, this implies $I$ is $2$-generated, so let us write $I=(r_{1},r_{2})D$. Now, $I]\in\text{ker}(\phi)$ implies $ID[\frac{1}{x}]$ is principal, so we write $ID[\frac{1}{x}]=(r_{1},r_{2})D[\frac{1}{x}]=rD[\frac{1}{x}]$. Note, we may assume without loss of generality that $r\in D$ because for any $\frac{a}{x^{n}}\in D[\frac{1}{x}]$, $\frac{a}{x^{n}}D[\frac{1}{x}]=aD[\frac{1}{x}]$. Now, the equality implies there exist $\frac{s_{1}}{x^{n_{1}}},\frac{s_{2}}{x^{n_{2}}}\in D[\frac{1}{x}]$ such that $r\frac{s_{1}}{x^{n_{1}}}=r_{1}$ and $r\frac{s_{2}}{x^{n_{2}}}=r_{2}$. Also, we must have $\frac{t_{1}}{x^{m_{1}}},\frac{t_{2}}{x^{m_{2}}}\in D[\frac{1}{x}]$ such that $r=r_{1}\frac{t_{1}}{x^{m_{1}}}+r_{2}\frac{t_{2}}{x^{m_{2}}}$. Without loss of generality, assume $n_{1}\geqslant n_{2}$ and $m_{1}\geqslant m_{2}$. Then, as $rs_{1}=r_{1}x^{n_{1}}$ and $rs_{2}=r_{2}x^{n_{2}}$, we see that $r$ divides $x^{n_{1}}r_{1}$ and $x^{n_{1}}r_{2}$ in $D$. Thus, $J=(\frac{x^{n_{1}}r_{1}}{r},\frac{x^{n_{1}}r_{2}}{r})D$ is an integral ideal of $D$, and $I\sim J$ as they differ by the principal ideal $(\frac{x^{n_{1}}}{r})D$. Now, from $r=r_{1}\frac{t_{1}}{x^{m_{1}}}+r_{2}\frac{t_{2}}{x^{m_{2}}}$ we get $1=t_{1}\frac{r_{1}}{rx^{m_{1}}}+t_{2}\frac{r_{1}}{rx_{m_{2}}}$, and multiplying by $x^{n_{1}+m_{1}}$, we get $x^{n_{1}+m_{1}}=t_{1}\frac{x^{n_{1}}r_{1}}{r}+t_{2}x^{m_{1}-m_{2}}\frac{x^{n}_{1}r_{2}}{r}\in J$.
The $\mathfrak{p}_{i}$ which divide $(x)$ are precisely those prime ideals which contain $x$, so we must have $J=\mathfrak{p}_{1}^{a_{1}}\mathfrak{p}_{2}^{a_{2}}\dotsm\mathfrak{p}_{k}^{a_{k}}$ with the possibility that $a_{i}=0$. Therefore, $I]=[J]=[\mathfrak{p}_{1}]^{a_{1}}[\mathfrak{p}_{2}]^{a_{2}}\dotsm[\mathfrak{p}_{k}]^{a_{k}}\in\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>, so the result follows from the first isomorphism theorem.
Notably, given $Cl(D)$ is torsion, Theorem 5.1 implies that any homomorphic image of $Cl(D)$ can be realized as the class group of such a localization of $D$. Claborn gives a similar, slightly weaker result in [3] where, given a Dedekind domain $A$ with arbitrary class group $G$ and subgroup $H$, he constructs a Dedekind polynomial ring over $A$ with class group isomorphic to $G/H$. As he answered the inverse class group problem for Dedekind domains, Theorem 5.1 brings us a step closer for overrings of $\mathcal{O}_{K}$. In essence, $\mathcal{O}_{K}[\frac{1}{x}]$ is very nearly a ring of integers itself. In fact, it is the precise analog if the base ring $\mathbb{Z}$ were replaced by $\mathbb{Z}[\frac{1}{x}]$ for some nonzero $x$ in the number field $K$. Note also that $Cl(\mathcal{O}_{K}[\frac{1}{x}])$ is finite and retains the useful property that each class contains a prime ideal. This follows directly from the correspondence between prime ideals of $\mathcal{O}_{K}[\frac{1}{x}]$ and prime ideals of $\mathcal{O}_{K}$ disjoint from $\{x^{n},\,n\in\mathbb{N}\}$. This allows for the application of many theorems on rings of integers to be applied to $\mathcal{O}_{K}[\frac{1}{x}]$. The following gives one such example.
Example 5.2.
Consider the ring of integers $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-14}]$ with class group $Cl_{K}\cong\mathbb{Z}/4\mathbb{Z}$ generated by $\mathfrak{p}=(3,1-\sqrt{-14})$. Writing $\mathfrak{q}=(2,\sqrt{-14})$, it is not difficult to show $\mathfrak{q}]=[\mathfrak{p}^{2}]$. Now $\mathfrak{q}^{2}=(2)$, so by Theorem 5.1 we have $Cl(\mathcal{O}_{K}[\frac{1}{2}])\cong\mathbb{Z}/2\mathbb{Z}$ implying $\mathcal{O}_{K}[\frac{1}{2}]=\mathbb{Z}[\sqrt{-14},\frac{1}{2}]$ is an HFD which is not a UFD.
Most significantly, Theorem 5.1 allows us to excise the minimal amount of $Cl_{K}$ and thus produce relatively large homomorphic images. In fact, we can realize any homomorphic image of $Cl_{K}$ in this way. In general, if $\gamma:Cl_{K}\rightarrow H$ is a homomorphism with $\text{ker}(\gamma)=\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{n}]\right$>, $Cl(\mathcal{O}_{K}[\frac{1}{\alpha}])\cong H$ where $(\alpha)=(\mathfrak{p}_{1}\mathfrak{p}_{2}\dotsm\mathfrak{p}_{n})^{h_{K}}$.
The remainder of this section will be dedicated to leveraging this fact to extend the techniques and theorems developed in sections 2 and 3. Our hope is that the Galois group of $K$ will act on the class group of $\mathcal{O}_{K}[\frac{1}{x}]$ in the same manner as $Cl_{K}$ which would allow us to place the same restrictions on the homomorphic images of the class group—further constraining the structure of $Cl_{K}$. Unfortunately, the action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in G$ and $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{x}])$ will not be well-defined in general as $\sigma(\frac{1}{x})=\sigma(x)^{-1}$ need not be in $\mathcal{O}_{K}[\frac{1}{x}]$ in general. However, in the case that $x\in\mathbb{Z}$, we avoid this issue as we will have $\sigma(\frac{1}{x})=\frac{1}{x}$ for all $\sigma\in G$.
Corollary 5.3.
Let $K$ be a Galois number field and $A$ the subgroup of $\text{Aut}(Cl_{K})$ induced by the action of of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$. Suppose that $x\in\mathcal{O}_{K}$ is a nonzero nonunit and
$(x)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}.$
Then $Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])\cong Cl_{K}/A\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> where $N(x)$ is the norm of $x$, and $A\left<X\right$> denotes the subgroup generated by the $A$-automorphic images of the elements in $X$.
Proof.
Recall that the ideal norm agrees with the element norm for principal ideals. Thus, as $K$ is Galois, we have $(N(x))=N((x))=\prod_{\sigma\in G}\sigma((x))=\prod_{\sigma\in G}\sigma(\mathfrak{p}_{1})^{n_{1}}\dotsm\sigma(\mathfrak{p}_{k})^{n_{k}}$. The theorem then follows by applying Theorem 5.1 to $N(x)$. ∎
Now, the group action of $G$ on the class group of $\mathcal{O}_{K}[\frac{1}{N(x)}]$ will be well-defined. More than this, it will be a Galois action.
Theorem 5.4.
For a Galois number field $K$ and $x\in\mathcal{O}_{K}$ a nonzero nonunit, the group action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in\text{Gal}(F/\mathbb{Q}),[\mathfrak{P}]\in Cl\left(\mathcal{O}_{K}[\frac{1}{N(x)}]\right)$ is a Galois action.
Proof.
It is sufficiently clear that conditions 1-3 from Definition 1.1 hold. Now, let $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ where $\mathfrak{P}$ is a prime ideal representative. If we write $\mathfrak{p}=\mathfrak{P}\mathbin{\cap}\mathcal{O}_{K}$, we have $\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{N(x)}]=\prod_{\sigma\in G}\sigma(\mathfrak{P})$, so the norm property (4) is also inherited from our original group action. ∎
Now, any restrictions the structure of $G$ places on $Cl_{K}$ must also hold for $Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ for any nonzero, nonunit $x$. In particular, if the set $\{[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\}$ is closed under the action of $G$, then we may apply our previous methods to $Cl_{K}/\left<[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\right$>. This is just another example of how misleading the quadratic case can be. As $\text{Gal}(K/\mathbb{Q})$ induces the subgroup $\{id,-id\}\leqslant\text{Aut}(Cl_{K})$ where $id$ is the identity and $id([I])=[I]^{-1}$, we see that any homomorphic image of $Cl_{K}$ can be realized by localizing at an integer—allowing significant restriction on $Cl_{K}$.
In general, it will be sufficient for a subgroup to be $\text{Aut}(Cl_{K})$-invariant. For example, if $G$ cannot admit a Galois action on $\mathbb{Z}/n\mathbb{Z}$, then we cannot have $Cl_{K}\cong\mathbb{Z}/an\mathbb{Z}$ for any $a\in\mathbb{N}$. This follows from the fact that $\mathbb{Z}/an\mathbb{Z}$ has a unique subgroup of order $n$. With this in hand, Theorem 3.4 becomes an immediate corollary of Theorem 3.1 in [6]. Furthermore, we can quickly improve upon Theorem 4.1.
Corollary 5.5.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, $Cl_{K}\not\cong\mathbb{Z}/np^{2}\mathbb{Z}$ for any $n\in\mathbb{N}$.
Once again, as we must have $h_{K}\equiv 0$ or $1$ mod $p$, this constitutes a serious restriction for extensions of (odd) prime degree. Now, note also that any Sylow $q$-subgroup of $Cl_{K}$ is invariant under automorphism. Thus, by the same logic, any restriction placed on $Cl_{K}$ may also be applied to its Sylow $q$-subgroups. Applying this also to Theorem 4.1, the following stronger result which appears in [11] is immediate.
Corollary 5.6.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, the Sylow $p$-subgroup of $Cl_{K}$ is not isomorphic to $\mathbb{Z}/p^{n}\mathbb{Z}$ for any $n\geqslant 2$.
In the same way, we can greatly strengthen Theorem 3.2.
Corollary 5.7.
Let $K$ be a Galois number field of degree $p^{r}$ and $S(q)$ a non-trivial Sylow $q$-subgroup of $Cl_{K}$. Then, $p=q$ or $|S(q)|\equiv 1$ mod $p$. Therefore, if $h_{k}=p^{r}q_{1}^{n_{1}}\dotsm q_{k}^{n_{k}}$, then $q_{i}^{n_{i}}\equiv 1$ mod $p$ for all $1\leqslant i\leqslant k$.
This was proven for abelian number fields by Fröhlich ([8]) and is a notably more restrictive condition. For example, observe that while Theorem 3.2 would not preclude an extension of degree $K:\mathbb{Q}]=3^{r}$ from having $h_{k}=55$, Corollary 5.7 shows this is impossible. In fact, such an extension cannot have $h_{K}=5n$ for any $n\in\mathbb{N}$ not divisible by $5$.
We conclude with a Jordan-Hölder like result which follows from Theorem 5.1.
Theorem 5.8.
Let $D$ be a Dedekind domain with finite class group $Cl(D)$ such that every class contains a prime ideal and $Cl(D)$. Let $R$ be an overring of $D$ which is minimal with respect to being a PID. Then there exists a finite sequence of adjacent domains $D:=D_{0}\subseteq D_{1}\subseteq D_{2}\subseteq\dotsm\subseteq D_{r}:=R$.
Proof.
Because $Cl(D)$ is abelian, we may choose an arbitrary element $\mathfrak{q}]\in Cl(D)$ of prime order $p$ where $\mathfrak{q}$ is a prime ideal representative of the class. Allowing $(a)=\mathfrak{q}^{p}$, by Theorem 5.1, $D_{1}:=D[\frac{1}{a}]$ has class group isomorphic to $Cl(D)/\left<[\mathfrak{q}]\right$>. We now show that $D\subseteq D[\frac{1}{a}]$ is a minimal ring extension.
Assume there exists some intermediate ring $D\subsetneq A\subseteq D[\frac{1}{a}]$. It is well known (see [9]) that
$A=\underset{\mathfrak{p}A\neq A}{\bigcap}D_{\mathfrak{p}}$
where the $\mathfrak{p}$ range over $\text{Spec}(D)$. Similarly,
$D=\underset{\mathfrak{m}\in\text{MaxSpec}(D)}{\bigcap}D_{\mathfrak{m}}=\underset{\mathfrak{p}\in\text{Spec}(D)}{\bigcap}D_{\mathfrak{p}}$
with the second equality due to $D$ being $1$-dimensional. As $A\neq D$, the above tells us there must exist some prime ideal $\mathfrak{p}^{\prime}\subseteq D$ such that $\mathfrak{p}^{\prime}A=A$ which implies some $d\in D$ becomes a unit in $A\subseteq D[\frac{1}{a}]$. Now, as $\mathfrak{p}]^{q}$ forms a minimal 0-sequence, we must have $a$ irreducible in $D$. Thus the saturation of $\{a^{n}\,|\,n\in\mathbb{N}\}$ in $D$ is simply $\{a^{n}\,|\,n\in\mathbb{N}\}$. Hence, $\frac{1}{a^{k}}\in A$ for some $k\in\mathbb{N}$ which implies $\frac{1}{a}\in A$, so we must have $A=D[\frac{1}{a}]$ as desired. Finally, as $Cl(D)$ is finite, we know that after a finite number of steps, we must have $Cl(D_{r})=1$, and thus $D_{r}$ is a PID.
Porism 5.9.
Let $D$ be a Dedekind domain with torsion class group and $a\in D$ irreducible. Then $D\subseteq D[\frac{1}{a}]$ is a minimal ring extension.
Example 5.10.
Recall $Cl_{K}\cong\mathbb{Z}/4\mathbb{Z}$ for $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-14}]$. Via the process described in the proof of Theorem 5.8, $\mathbb{Z}[\sqrt{-14}]\subseteq\mathbb{Z}[\sqrt{-14},\frac{1}{2}]\subseteq\mathbb{Z}[\sqrt{-14},\frac{1}{6}]$ is a series of adjacent domains with $\mathbb{Z}[\sqrt{-14},\frac{1}{6}]$ a PID.
Thus, we see that any ring of integers, in a sense, is finitely many steps away from being a PID.