### v2

# Galois Action and Localization in Number Fields

Abstract.
For a Galois number field $K$, the Galois group $\text{Gal}(K/\mathbb{Q})$ acts on the class group $Cl_{K}$ in a very natural way: $\sigma\cdot[I]=[\sigma(I)]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$, $I]\in Cl_{K}$. In this paper, we will explore how the unique properties of this group action work together to elucidate the relationship between these two groups. While previous work on this problem has focused on representation theory, we take a direct approach to some classical and new problems. The paper concludes with an exploration of the class groups of overrings of the ring of integers $\mathcal{O}_{K}$. These turn out to be powerful tools for understanding $Cl_{K}$ and overrings of $\mathcal{O}_{K}$.
1. Introduction and Notation
Throughout this paper $K$ will denote a Galois number field with Galois group $G:=\text{Gal}(K/\mathbb{Q})$, and $\mathcal{O}_{K}$ its ring of integers with class group $Cl_{K}$ and class number $h_{K}=|Cl_{K}|$. The Galois group of $K$ acts on the class group in a very natural way. For any $\sigma\in G$ and $I]\in Cl_{K}$, we can define the action $\sigma\cdot[I]=[\sigma(I)]$. This is a well-defined group action which provides us with some tools for characterization the relationship between the two groups. More than this, we also have the peculiar property that $\sigma\cdot([I][J])=(\sigma\cdot[I])(\sigma\cdot[J])$ for any $\sigma\in G$ and $I],[J]\in Cl_{K}$. This induces a map from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\mapsto\bar{\sigma}$ where $\bar{\sigma}([I])=[\sigma(I)]$. Finally, we have the “norm property” that $\prod_{\sigma\in G}\sigma\cdot[I]=\left[\prod_{\sigma\in G}\sigma(I)\right]=\text{Prin}(\mathcal{O}_{K})$ for any $I]\in Cl_{K}$—for details, see [13]. Note, in this paper we will study the case where $K$ is Galois over $\mathbb{Q}$ with the knowledge that the theorems herein apply also to the case when $K/L$ is a Galois extension of number fields with $\mathcal{O}_{L}$ a principal ideal domain (PID). Now, these characteristics of the action motivate the following definition.
Definition 1.1.
Let $G$ and $A$ be groups with $A$ abelian, and let
$\alpha:G\times A\longrightarrow A,\quad(g,a)\mapsto g\cdot a$
be a map. If $\alpha$ satisfies the following properties:
  1. (1)
    $g_{1}\cdot(g_{2}\cdot a)=(g_{1}g_{2})\cdot a$ for all $g_{1},g_{2}\in G$, $a\in A$,
  2. (2)
    $e_{G}\cdot a=a$ for all $a\in A$,
  3. (3)
    $g\cdot(a_{1}a_{2})=(g\cdot a_{1})(g\cdot a_{2})$ for all $g\in G$, $a_{1},a_{2}\in A$,
  4. (4)
    $\prod_{g\in G}(g\cdot a)=e_{A}$ for all $a\in A$,
then we say that $\alpha$ is a norm-like action.
Throughout this paper, we will see how properties 1-4 can be used to place strong restrictions on the structure of $Cl_{K}$ given $G$ and vice versa. We will place a particular emphasis on techniques developed and some applications to factorization theory.
The previous work done on this subject has been primarily through the lens of representation theory—considering $Cl_{K}$ as a $G$-module. This approach was utilized by Fröhlich in 1952 ([9]) and subsequently many others. Cornell and Rosen used this approach in [6] to give results on the structure of $Cl_{K}$ when $h_{K}$ is known. Others such as Lemmermeyer ([12]) and Iwasawa ([11]) used the $G$-module structure of $Cl_{K}$ along with assumptions on intermediate fields to give various results on the $p$-rank of the class group. For a thorough review of the previous work done in this area, see [14].
In the following section, we will take a closer look at the conditions of Definition 1.1 and their implications for norm-like actions in general. In Section 3, we explore how conditions 1-2 together with 4 place restrictions of the structure of $Cl_{K}$ for a Galois number field of given degree. In Section 4, we focus on applications of condition 3 to the inverse class group problem. In particular, which number fields $K$ can have a given abelian group as their class group. In Section 5, the authors determine the structure of the class group for localizations of the form $\mathcal{O}_{K}[\frac{1}{x}]$ which ultimately strengthens many of the previously developed results and leads to some interesting observations about rings of integers. Finally, Section 6 concludes with a more general exploration of overrings of Dedekind domains, culminating with a characterization of those overrings whose class groups admit a norm-like group action.
2. Norm-Like Group Actions
In this section, we will consider norm-like group actions directly. The results and observation made here will be used throughout the remainder of the paper. First, let us recall the definitions of a norm-like action.
Definition 1.1.
Let $G$ and $A$ be groups with $A$ abelian, and let
$\alpha:G\times A\longrightarrow A,\quad(g,a)\mapsto g\cdot a$
be a map. If $\alpha$ satisfies the following properties:
  1. (1)
    $g_{1}\cdot(g_{2}\cdot a)=(g_{1}g_{2})\cdot a$ for all $g_{1},g_{2}\in G$, $a\in A$,
  2. (2)
    $e_{G}\cdot a=a$ for all $a\in A$,
  3. (3)
    $g\cdot(a_{1}a_{2})=(g\cdot a_{1})(g\cdot a_{2})$ for all $g\in G$, $a_{1},a_{2}\in A$,
  4. (4)
    $\prod_{g\in G}(g\cdot a)=e_{A}$ for all $a\in A$,
then we say that $\alpha$ is a norm-like action.
Conditions $1$ and $2$ are those of a typical group action. The definition diverges with condition $3$ which ensures the action of each element $g$ on $A$ behaves like a homomorphism. In fact, each element of $G$ can be identified with an endomorphism of $A$ via the map $g\mapsto\phi$ where $\phi(a)=g\cdot a$. If $A$ is finite, in the case where $g\cdot a=e_{A}\implies a=e_{A}$, this map defines a group homomorphism from $G$ to $\text{Aut}(A)$, the kernel of which is the set $\{g\in G\,|\,g\cdot a=a\,\forall\,a\in A\}$. Note, this is precisely the kernel of the group action. The implications of this are closely considered in Section 4. We also note that condition $3$ guarantees $g\cdot e_{A}=e_{A}$ for all $g\in G$. Thus, a norm-like group action will be transitive if and only if $A$ is trivial.
Let us consider a group action of $G$ on $A$ satisfying conditions 131-3. First, it suffices to check condition $4$ for the generators of $A$. Now, consider the subgroup $H:=\{\prod_{g\in G}g\cdot a\,|\,a\in A\}\unlhd A$. For any $x\in H$, we have $x=\prod_{g\in G}g\cdot a^{\prime}$ for some $a^{\prime}\in A$. Thus, for any $g^{\prime}\in G$, $g^{\prime}\cdot x=g^{\prime}\cdot\left(\prod_{g\in G}g\cdot a^{\prime}\right)=\prod_{g\in G}(g^{\prime}*g)\cdot a^{\prime}=\prod_{g\in G}g\cdot a^{\prime}=x$. Hence, $H$ is a subgroup of the elements of $A$ with trivial orbit. Therefore, if $e_{A}$ is the only element with a trivial orbit, condition $4$ follows from 131-3.
Now, let $A$ be an arbitrary abelian group, and $G=\{id,-id\}\leqslant\text{Aut}(A)$ where $id$ is the identity, and id(a)=a-id(a)=-a for all $a\in A$ which is a automorphism just in case $A$ is abelian. It is easy to verify that the natural action $g\cdot a=g(a)$ is norm-like. This case has been of particular interest over the last three years as the study of weighted zero sum sequences has become popular in connection with normset factorization (see [1] and [8]). In particular, $\{id,-id\}$ is the subgroup induced by the action of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ when $K$ is quadratic. As we have seen, $|G|=2$ places no restriction on $A$. As we will see throughout this paper, this is a very unique case. Even the $|G|=3$ case is much more restrictive and the relevant arithmetic far more complicated.
Example 2.1.
Cubic Example: $G\cong\mathbb{Z}/3\mathbb{Z}$ acting on $A\cong\mathbb{Z}/7\mathbb{Z}$ Also, any of the subgroups of $SL_{2}(\mathbb{Z}/2\mathbb{Z})$ (order $6$) acting on $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ seem to work. There might be something here. Probably want a non-example in the cubic case as well.Is this unique in the sense that the action of all subgroups of $\text{Aut}(A)$ work via the natural group action.
3. Direct Consequences of the Group Action
In this section, we produce results on the structure of the class group of Galois number fields which follow directly from the action of the Galois group $G=\text{Gal}(F/\mathbb{Q})$ on the class group $Cl_{K}$ in conjunction with the norm property $\prod_{\sigma\in G}\sigma(I)\in\text{Prin}(\mathcal{O}_{K})$ for any ideal $I\subseteq\mathcal{O}_{K}$. We begin with a lemma which we will use extensively throughout this paper.
Lemma 3.1.
Let $K$ be a Galois number field with $G=\text{Gal}(K/\mathbb{Q})$ and class group $Cl_{K}$. If $G$ acts trivially on $I]\in Cl_{K}$, the order of $[I]$ divides $|G|=[K:\mathbb{Q}]$.
Proof.
Assume $G$ acts trivially on $I]\in Cl_{K}$. Now, by the norm property, we have
$\prod_{\sigma\in G}\sigma(I)\in\text{Prin}(\mathcal{O}_{K}).$
Thus, we have
$\prod_{\sigma\in G}[\sigma(I)]=\text{Prin}(\mathcal{O}_{K})\Rightarrow{}\prod_{\sigma\in G}\sigma\cdot[I]=\text{Prin}(\mathcal{O}_{K}).$
As $G$ acts trivially on $[I]$, we have
$\text{Prin}(\mathcal{O}_{K})=\prod_{\sigma\in G}\sigma\cdot[I]=\prod_{\sigma\in G}[I]=[I]^{|G|}.$
This completes the proof. ∎
The first result we present follows from Lemma 2 in [6] using the representation theoretic approach. We will prove it directly here and generalize the result in section $4$ with the techniques developed therein.
Theorem 3.2.
Let $K$ be a Galois number field of degree $p^{r}$ where $p$ is prime. Then, $h_{K}\equiv 0\text{ or }1\text{ mod }p$.
Proof.
Assume that $h_{K}\not\equiv 1$ mod $p$. Once again, let $G$ act on $Cl_{K}$ via the action $\sigma\cdot[I]=[\sigma(I)]$. By the orbit-stabilizer theorem, the lengths of the orbits of this group action divide $|G|=[K:\mathbb{Q}]=p^{r}$. In particular, they must be elements of the set $\{1,p,p^{2},\dots,p^{r}\}$.
Now, the identity of $Cl_{K}$ has a trivial orbit of order $1$. The orbits partition $Cl_{K}$, and all non-trivial orbits have order congruent to 0 mod $p$. Hence, as we assumed $h_{K}\not\equiv 1$ mod $p$, there must be some non-identity class of $Cl_{K}$ in a trivial orbit. Call this element $J]\neq\text{Prin}(\mathcal{O}_{K})$. By Lemma 3.1, $[J]$ has order dividing $p^{r}$, and as we assumed $[J]$ was not the identity, $p$ must divide the order of $[J]$. Therefore, by Lagrange’s theorem, $p\text{ divides }h_{K}$, so $h_{K}\equiv 0$ mod $p$. This completes the proof.
This theorem demonstrates one of the many ways in which quadratic number fields are exceptional. Beyond their peculiarity of being Galois in general, Theorem 3.2 leaves all class numbers permissible in the quadratic case. The same is true of the following.
Theorem 3.3.
If $K$ is a Galois number field of degree $n$ and $p$ the smallest prime divisor of $n$, $h_{k}=1$ or $h_{k}\geqslant p$.
Proof.
Assume for the purpose of contradiction that $1<h_{k}<p$. As $h_{k}<p$, the smallest prime divisor of $|G|$, the orbit-stabilizer theorem tells us every element of $Cl_{K}$ has a trivial orbit under the action of $G$ on $Cl_{K}$. As $h_{k}>1$, we may take $J]\neq\text{Prin}(\mathcal{O}_{K})$ in $Cl_{K}$. Then, by Lemma 3.1, we must have $J]^{|G|}=\text{Prin}(\mathcal{O}_{K})$, so the order of $[J]$ divides $|G|=n$. However, this is a contradiction as $1<|[J]|<h_{k}<p$, and $p$ is the smallest prime divisor of $n$. ∎
The question of whether every finite abelian group can be realized as the class group of a quadratic ring of integers is a famous open problem. We will consider some related questions in the next section. For now, we now turn our attention to some results on Galois number fields of odd degree.
Theorem 3.4.
Let $K$ be a Galois number field of odd degree. Then, $Cl_{K}\ncong\mathbb{Z}/2n\mathbb{Z}$ for any $n\in\mathbb{N}$.
Proof.
Assume for the purpose of contradiction that $Cl_{K}\cong\mathbb{Z}/2n\mathbb{Z}$. Then, $Cl_{K}$ has a unique element of order 2, call it $[J]$. Thus, as each $\sigma\in G$ is an automorphism, we must have $\sigma\cdot[J]=[\sigma(J)]=[J]$ for all $\sigma\in G$. Hence, $G$ acts trivially on $[J]$, so by Lemma 3.1 the order of $[J]$ divides $|G|$. However, this is a contradiction as $[J]$ has order $2$ and $|G|$ is odd. ∎
In particular, for a Galois number field of odd degree, $Cl_{K}$ cannot have a unique invariant factor of even order. Notably, this precludes cyclic groups of even order. Now, from Theorem 3.4 we get the following factorization result which first appears as Lemma 3.5 in [7].
Corollary 3.5.
Let $K$ be a Galois number field of odd degree. Then $\mathcal{O}_{K}$ is an HFD if and only if $\mathcal{O}_{K}$ is a UFD.
Proof.
A well-known result from Carlitz ([2]) tells us $\mathcal{O}_{K}$ is an HFD if and only if $h_{K}=1$ or $2$. Noting that Theorem 3.4 disallows $h_{K}=2$ and $\mathcal{O}_{K}$ is a UFD if and only if $h_{K}=1$ completes the proof. ∎
Note that this result also follows directly from Theorem 3.3. We conclude this section with a specific application of these results to factorization in cyclotomic number fields. These number fields in particular played a central role in the development of modern factorization theory. In his attempt to prove Fermat’s Last Theorem, Ernst Kummer erroneously assumed that the cyclotomic ring of integers $\mathbb{Z}[\xi_{p}]$ was a unique factorization domain for any primitive $p^{th}$ root of unity. It was Dedekind’s remarkable observation that while unique factorization for elements was lost, unique factorization of ideals was retained for all rings of integers. Thus, we are able to turn questions about the factorization of elements in rings of integers into questions about ideals and, in turn, the class group. Being Galois in general, the cyclotomic number fields provide an apt setting for the applications of the results we have developed in this section.
Theorem 3.6.
Let $p<23$ be an odd prime and $a\in\mathbb{Z}$ not divisible by the $p^{th}$ power of any prime. If $K$ is the splitting field of $x^{p}-a$, $\mathcal{O}_{K}$ is an HFD if and only if it is $\mathcal{O}_{K}$ is a UFD.
Proof.
We first note that, as the splitting field of $x^{p}-a$, $K$ is Galois over $\mathbb{Q}$. Furthermore, we have $K\cong\mathbb{Q}(\omega,\alpha)$ where $\omega$ is a primitive $p^{th}$ root of unity and $\alpha$ a root of $x^{p}-a$. This gives us the following lattice.
$\mathbb{Q}$$\mathbb{Q}(\omega)$$\mathbb{Q}(\alpha)$$\mathbb{Q}(\omega,\alpha)$
Where $\mathbb{Q}(\omega,\alpha)$ is Galois over $\mathbb{Q}(\omega)$ of degree $p$. Now, as $p<23$, $\mathbb{Z}[\omega]$ is a UFD, so the result follows from Corollary 3.5.
[…]
Recall that the action of $G=\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ induces a homomorphism from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\overset{\psi}{\mapsto}\phi$ where $\phi([I])=\sigma\cdot[I]=[\sigma(I)]$. In this section, we will see how this map can be used to further restrict the structure of $Cl_{K}$ for a Galois number field $K$. Conversely, we will use this tool to help answer questions related to the inverse class group problem. This terminology is used in the spirit of the inverse Galois problem. In particular, the inverse class group problem asks if an arbitrary abelian group can be realized as the class group of a ring of integers. This question was famously answered in the positive for the more general class of Dedekind domains by Claborn in [5]. In the case of imaginary quadratics, the answer is known to be no. Certain groups can be shown by brute force not to appear, the smallest among them being $(\mathbb{Z}/3\mathbb{Z})^{3}$. The question remains open for rings of integers in general. In this section, for a given finite abelian group $A$, we will explore which rings of integers may admit $Cl_{K}\cong A$. We begin with a direct proof of another result from [6] which we will return to with some new found tools in section 4.
Theorem 4.1.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, $Cl_{K}\ncong\mathbb{Z}/p^{n}\mathbb{Z}$ for $n\geqslant 2$.
Proof.
[…]
Assume $H$ is trivial. Then, each element of $G$ acts trivially on the class group. Therefore, by Lemma 3.1, every element of $Cl_{K}$ has order dividing $|G|=p$, but this contradicts our assumption that $Cl_{K}\cong\mathbb{Z}/p^{n}\mathbb{Z}$ with $n\geqslant 2$. Hence, $H$ is the unique subgroup of order $p$.
[…]
$\prod_{\sigma\in G}\sigma(J)\in\text{Prin}(\mathcal{O}_{K})$
implies
$\prod_{\sigma\in G}[\sigma(J)]=\prod_{\sigma\in G}\sigma[(J)]=\prod_{i=0}^{p-1}\psi^{i}([J])=\text{Prin}(\mathcal{O}_{K})$
Furthermore, $\psi^{i}([J])=(\gamma^{(p-1)p^{n-2}})^{i}([J])=[J]^{\left(\alpha^{(p-1)p^{n-2}}\right)^{i}}$. Note, $(p-1)p^{n-2}=\varphi(p^{n-1})$, so $\alpha^{(p-1)p^{n-2}}\equiv 1$ mod $p^{n-1}$. Also, because $\alpha$ is a primitive root mod $p^{n}$, each $(\alpha^{(p-1)p^{n-2}})^{i}=\alpha^{i(p-1)p^{n-2}}$ is unique mod $p^{n}$ for each $0\leqslant i\leqslant p-1$. Hence, they are precisely the elements $\{1,p^{n-1}+1,2p^{n-1}+1,\dots,(p-1)p^{n-1}+1\}$ mod $p^{n}$. Taking their sum, we get
$\sum_{i=0}^{p-1}\left(1+i\cdot p^{n-1}\right)=p+p^{n-1}\sum_{i=1}^{p-1}i=p+p^{n-1}\cdot\frac{(p-1)p}{2}=p+\frac{p-1}{2}\cdot p^{n}\equiv p\text{ mod }p^{n}$
Hence,
$\prod_{i=0}^{p-1}\psi^{i}([J])=\prod_{i=0}^{p-1}[J]^{\left(\alpha^{(p-1)p^{n-2}}\right)^{i}}=[J]^{\sum_{i=0}^{p-1}1+i\cdot p^{n-1}}=[J]^{p}=\text{Prin}(\mathcal{O}_{K})$
But this implies the order of $[J]$ divides $p$, and $[J]$ is of order $p^{n}$ with $n\geqslant 2$. This is a contradiction. Therefore, $Cl_{K}$ cannot be $\mathbb{Z}/p^{n}\mathbb{Z}$ for any $n\geqslant 2$.
Recall that Theorem 3.2 tells us (in particular) that for a Galois number field of degree $p$, we must have $h_{k}\equiv 0$ or $1$ mod $p$. Thus, Theorem 4.1 gives us a further restriction in the first case on the structure of the class group. We now shift our attention to the inverse class group. First, we consider the case when $h_{k}$ is assumed to be an odd prime rather than $K:\mathbb{Q}]$. Note, this is equivalent to asking which number fields $K$ can have $Cl_{K}\cong\mathbb{Z}/p\mathbb{Z}$.
Theorem 4.2.
Let $K$ a Galois number field with $n=[K:\mathbb{Q}]$ and $h_{K}=p$ prime. Then, $p|n$ or gcd(p1,n)>1(p-1,n)>1.
Proof.
[…]
This theorem tells us that having class group $\mathbb{Z}/p\mathbb{Z}$ is a relatively restrictive condition. For example, only Galois number fields of degree divisible by $2$ or $17$ can have class group $\mathbb{Z}/17\mathbb{Z}$. We continue this section with a few specific examples of the inverse class group problem which present new techniques and highlight the interesting interplay between the induced homomorphism and the norm property of the group action.
Example 4.3.
Let us assume $Cl_{K}\cong\mathbb{Z}/13\mathbb{Z}$ for some Galois number field $K$. None of the theorems developed thus far preclude a cubic Galois extension from having such a class group, and indeed we see that the number field with defining polynomial $x^{3}-x^{2}-354x-2441$ is one such example. Now, any cubic Galois extension has $G\cong\mathbb{Z}/3\mathbb{Z}$, so the induced homomorphism $\psi:G\rightarrow\text{Aut}(Cl_{K})$ must be injective as the alternative would imply there exists an element in $\mathbb{Z}/13\mathbb{Z}$ of order $|G|=3$.
[…]
The following theorem is inspired by this example and demonstrates that the phenomenon described at the end must occur in general for $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$.
Theorem 4.4.
Let $p$ be prime and $\phi=(a_{1},a_{2},...,a_{p-1})$ a cyclic generator of $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$. If $n$ properly divides p1p-1, then the elements in each disjoint cycle of $\phi^{n}$ sum to 0 mod $p$.
Proof.
Say $\phi(1)=a$, then we must have $\phi(x)=a\cdot x$ for any $x\in\mathbb{Z}/p\mathbb{Z}$. As $\phi$ generates $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$, this also implies that $a$ is a primitive root mod $p$. Now, any $x\in\mathbb{Z}/p\mathbb{Z}$ is contained in a (p1)/n(p-1)/n cycle in $\phi^{n}$. Thus, the cycle is of the form $(x\,\,\,xa^{n}\,\,\,xa^{2n}\dotsm xa^{(\frac{p-1}{n}-1)n})$. Taking the sum, we get
$x\sum_{r=0}^{\frac{p-1}{n}-1}(a^{n})^{r}=x\cdot\frac{1-(a^{n})^{\frac{p-1}{n}}}{1-a^{n}}=x\cdot\frac{1-a^{p-1}}{1-a^{n}}$
By Fermat’s Little Theorem, we have $1-a^{p-1}\equiv 0$ mod $p$. The result then follows from the fact that $n$ properly divides p1p-1, so $1-a^{n}\not\equiv 0$ mod $p$.
Example 4.5.
Consider a Galois number field $K$ with class group $Cl_{K}\cong(\mathbb{Z}/2\mathbb{Z})^{3}$. It is well known that $\text{Aut}((\mathbb{Z}/2\mathbb{Z})^{3})\cong\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$, the group of invertible $3\times 3$ matrices over the field of two elements with order $|\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})|=168=2^{3}\cdot 3\cdot 7$. Employing the same methods as in the proof of Theorem 4.2, we immediately see that the order of the extension $n=[K:\mathbb{Q}]$ must be divisible by $2,3$ or $7$. However, Theorem 3.2 tells us that no Galois extension of degree $3^{r}$ can have class group $(\mathbb{Z}/2\mathbb{Z})^{3}$ for any $r\geqslant 1$. We will now show that in fact no Galois number field of degree $3m$ where $\text{gcd}(2,m)=\text{gcd}(7,m)=1$ can have class group $(\mathbb{Z}/2\mathbb{Z})^{3}$.
Assume $G=3m$ where $m$ is not divisible by $2$ or $7$. Once again, the homomorphism $\psi:G\rightarrow\text{Aut}(Cl_{K})\cong\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ must be non-trivial as $Cl_{K}$ has no non-identity elements of order dividing $|G|=3m$. Thus, $im(\psi)$ must be a subgroup of order $3$ by the first isomorphism theorem. Now, consider the following element of order $3$ in $\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ and its powers:
$E=\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&1&1\end{bmatrix}\hskip 14.22636ptE^{2}=\begin{bmatrix}1&0&0\\ 0&1&1\\ 0&1&0\end{bmatrix}\hskip 14.22636ptE^{3}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
Now, $E$ has characteristic polynomial $p_{E}(x)=(x+1)(x^{2}+x+1)$, so we see it has eigenvalue $\lambda=1$. From here, one can show $(1,0,0)^{T}$ is a corresponding eigenvector. That is, $E$ fixes the vector $(1,0,0)^{T}$, and so $E^{2}$ must also. Furthermore, by the second Sylow theorem and the fact that similar matrices have the same eigenspaces, we conclude that any element of order $3$ in $\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ fixes $(1,0,0)^{T}$. Therefore, as $im(\psi)$ is a subgroup of order $3$, we must have $(1,0,0)\in(\mathbb{Z}/2\mathbb{Z})^{3}$ fixed by every element of $G$, but this implies the order of $(1,0,0)$, namely $2$, divides $|G|=3m$, a contradiction. Therefore, we see that some degrees such as $K:\mathbb{Q}]=15$ which were not previously disallowed are impossible for $Cl_{K}\cong(\mathbb{Z}/2\mathbb{Z})^{3}$.
Similar methods can be employed to investigate the inverse class group problem for $p$-elementary abelian groups as well as cyclic groups of order $2^{m}p^{r}$ with $m\in\{0,1\}$ and $r\geqslant 1$ where $p$ is an odd prime.
5. Localizations of Rings of Integers
In this section, we will use localization to strengthen some of the results developed previously. First, recall that any overring—in particular, any localization—of a Dedekind domain remains Dedekind. Intuitively, as a localization of $R$ is attained by turning a set of elements $S\subseteq R$ into units, we expect the class group of $R_{S}$ to be smaller than that of $R$ as those non-principal ideals $I\subseteq R$ for which $I\cap S\neq\emptyset$ will become principal. As we will demonstrate, “small” localizations will tend to lead to small changes in the class group. In the context of rings of algebraic integers, this will be useful in further restricting the possible structure of $Cl_{K}$.
Theorem 5.1.
Let $D$ be a Dedekind domain and $x\in D$ a nonzero nonunit with
$(x)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}$
Then $Cl(D[\frac{1}{x}])\cong Cl(D)/\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>.
Note, $D[\frac{1}{x}]$ is the localization of $D$ at the multiplicative set $\{x^{n},\,n\in\mathbb{N}\}$ which is the complement of the set-theoretic union of all prime ideals not containing $x$.
Proof.
Let $\phi:Cl(D)\rightarrow Cl(D[\frac{1}{x}])$ be the canonical map $I]\overset{\phi}{\mapsto}\left[ID[\frac{1}{x}]\right]$. First we will show that $\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> is contained in $\text{ker}(\phi)$. It suffices to show that $\mathfrak{p}_{i}$ becomes principal in $D[\frac{1}{x}]$ for any $1\leqslant i\leqslant k$. This is clearly the case as $x\in\mathfrak{p}_{i}$, so $\mathfrak{p}_{i}D[\frac{1}{x}]=D[\frac{1}{x}]$.
Now, let $I]\in\text{ker}(\phi)$. The case when $I]=\text{Prin}(D)$ is trivial, so we will assume $I$ is a non-principal, integral ideal. As $D$ is Dedekind, this implies $I$ is $2$-generated, so let us write $I=(r_{1},r_{2})D$. Now, $I]\in\text{ker}(\phi)$ implies $ID[\frac{1}{x}]$ is principal, so we write $ID[\frac{1}{x}]=(r_{1},r_{2})D[\frac{1}{x}]=rD[\frac{1}{x}]$. Note, we may assume without loss of generality that $r\in D$ because for any $\frac{a}{x^{n}}\in D[\frac{1}{x}]$, $\frac{a}{x^{n}}D[\frac{1}{x}]=aD[\frac{1}{x}]$. Now, the equality implies there exist $\frac{s_{1}}{x^{n_{1}}},\frac{s_{2}}{x^{n_{2}}}\in D[\frac{1}{x}]$ such that $r\frac{s_{1}}{x^{n_{1}}}=r_{1}$ and $r\frac{s_{2}}{x^{n_{2}}}=r_{2}$. Also, we must have $\frac{t_{1}}{x^{m_{1}}},\frac{t_{2}}{x^{m_{2}}}\in D[\frac{1}{x}]$ such that $r=r_{1}\frac{t_{1}}{x^{m_{1}}}+r_{2}\frac{t_{2}}{x^{m_{2}}}$. Without loss of generality, assume $n_{1}\geqslant n_{2}$ and $m_{1}\geqslant m_{2}$. Then, as $rs_{1}=r_{1}x^{n_{1}}$ and $rs_{2}=r_{2}x^{n_{2}}$, we see that $r$ divides $x^{n_{1}}r_{1}$ and $x^{n_{1}}r_{2}$ in $D$. Thus, $J=(\frac{x^{n_{1}}r_{1}}{r},\frac{x^{n_{1}}r_{2}}{r})D$ is an integral ideal of $D$, and $I\sim J$ as they differ by the principal ideal $(\frac{x^{n_{1}}}{r})D$. Now, from $r=r_{1}\frac{t_{1}}{x^{m_{1}}}+r_{2}\frac{t_{2}}{x^{m_{2}}}$ we get $1=t_{1}\frac{r_{1}}{rx^{m_{1}}}+t_{2}\frac{r_{1}}{rx_{m_{2}}}$, and multiplying by $x^{n_{1}+m_{1}}$, we get $x^{n_{1}+m_{1}}=t_{1}\frac{x^{n_{1}}r_{1}}{r}+t_{2}x^{m_{1}-m_{2}}\frac{x^{n}_{1}r_{2}}{r}\in J$.
The $\mathfrak{p}_{i}$ which divide $(x)$ are precisely those prime ideals which contain $x$, so we must have $J=\mathfrak{p}_{1}^{a_{1}}\mathfrak{p}_{2}^{a_{2}}\dotsm\mathfrak{p}_{k}^{a_{k}}$ with the possibility that $a_{i}=0$. Therefore, $I]=[J]=[\mathfrak{p}_{1}]^{a_{1}}[\mathfrak{p}_{2}]^{a_{2}}\dotsm[\mathfrak{p}_{k}]^{a_{k}}\in\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>, so the result follows from the first isomorphism theorem.
Notably, given $Cl(D)$ is torsion, Theorem 5.1 implies that any homomorphic image of $Cl(D)$ can be realized as the class group of such a localization of $D$. Claborn gives a similar, slightly weaker result in [4] where, given a Dedekind domain $A$ with arbitrary class group $G$ and subgroup $H$, he constructs a Dedekind polynomial ring over $A$ with class group isomorphic to $G/H$. As he answered the inverse class group problem for Dedekind domains, Theorem 5.1 brings us a step closer for overrings of $\mathcal{O}_{K}$. In essence, $\mathcal{O}_{K}[\frac{1}{x}]$ is very nearly a ring of integers itself. In fact, it is the precise analog if the base ring $\mathbb{Z}$ were replaced by $\mathbb{Z}[\frac{1}{x}]$ for some nonzero $x$ in the number field $K$. Note also that $Cl(\mathcal{O}_{K}[\frac{1}{x}])$ is finite and retains the useful property that each class contains a prime ideal. This follows directly from the correspondence between prime ideals of $\mathcal{O}_{K}[\frac{1}{x}]$ and prime ideals of $\mathcal{O}_{K}$ disjoint from $\{x^{n},\,n\in\mathbb{N}\}$. This allows for the application of many theorems on rings of integers to be applied to $\mathcal{O}_{K}[\frac{1}{x}]$. The following gives one such example.
Example 5.2.
Consider the ring of integers $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-14}]$ with class group $Cl_{K}\cong\mathbb{Z}/4\mathbb{Z}$ generated by $\mathfrak{p}=(3,1-\sqrt{-14})$. Writing $\mathfrak{q}=(2,\sqrt{-14})$, it is not difficult to show $\mathfrak{q}]=[\mathfrak{p}^{2}]$. Now $\mathfrak{q}^{2}=(2)$, so by Theorem 5.1 we have $Cl(\mathcal{O}_{K}[\frac{1}{2}])\cong\mathbb{Z}/2\mathbb{Z}$ implying $\mathcal{O}_{K}[\frac{1}{2}]=\mathbb{Z}[\sqrt{-14},\frac{1}{2}]$ is an HFD which is not a UFD.
Most significantly, Theorem 5.1 allows us to excise the minimal amount of $Cl_{K}$ and thus produce relatively large homomorphic images. In fact, we can realize any homomorphic image of $Cl_{K}$ in this way. In general, if $\gamma:Cl_{K}\rightarrow H$ is a homomorphism with $\text{ker}(\gamma)=\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{n}]\right$>, $Cl(\mathcal{O}_{K}[\frac{1}{\alpha}])\cong H$ where $(\alpha)=(\mathfrak{p}_{1}\mathfrak{p}_{2}\dotsm\mathfrak{p}_{n})^{h_{K}}$.
The remainder of this section will be dedicated to leveraging this fact to extend the techniques and theorems developed in sections 2 and 3. Our hope is that the Galois group of $K$ will act on the class group of $\mathcal{O}_{K}[\frac{1}{x}]$ in the same manner as $\text{Cl}_{K}$ which would allow us to place the same restrictions on the homomorphic images of the class group—further constraining the structure of $Cl_{K}$. Unfortunately, the action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in G$ and $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{x}])$ will not be well-defined in general as $\sigma(\frac{1}{x})=\sigma(x)^{-1}$ need not be in $\mathcal{O}_{K}[\frac{1}{x}]$ in general. However, in the case that $x\in\mathbb{Z}$, we avoid this issue as we will have $\sigma(\frac{1}{x})=\frac{1}{x}$ for all $\sigma\in G$.
Corollary 5.3.
Let $K$ be a Galois number field and $A$ the subgroup of $\text{Aut}(Cl_{K})$ induced by the action of of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$. Suppose that $x\in\mathcal{O}_{K}$ is a nonzero nonunit and
$(x)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}.$
Then $Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])\cong Cl_{K}/A\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> where $N(x)$ is the norm of $x$, and $A\left<X\right$> denotes the subgroup generated by the $A$-automorphic images of the elements in $X$.
Proof.
Recall that the ideal norm agrees with the element norm for principal ideals. Thus, as $K$ is Galois, we have $(N(x))=N((x))=\prod_{\sigma\in G}\sigma((x))=\prod_{\sigma\in G}\sigma(\mathfrak{p}_{1})^{n_{1}}\dotsm\sigma(\mathfrak{p}_{k})^{n_{k}}$. The theorem then follows by applying Theorem 5.1 to $N(x)$. ∎
Now, the group action of $G$ on the class group of $\mathcal{O}_{K}[\frac{1}{N(x)}]$ will be well-defined. More than this, it will be a Galois action.
Theorem 5.4.
For a Galois number field $K$ and $x\in\mathcal{O}_{K}$ a nonzero nonunit, the group action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in\text{Gal}(F/\mathbb{Q}),[\mathfrak{P}]\in Cl\left(\mathcal{O}_{K}[\frac{1}{N(x)}]\right)$ is a norm-like action.
Proof.
It is sufficiently clear that conditions 1-3 from Definition 1.1 hold. Now, let $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ where $\mathfrak{P}$ is a prime ideal representative. If we write $\mathfrak{p}=\mathfrak{P}\cap\mathcal{O}_{K}$, we have $\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{N(x)}]=\prod_{\sigma\in G}\sigma(\mathfrak{P})$, so the norm property (4) is also inherited from our original group action. ∎
Now, any restrictions the structure of $G$ places on $Cl_{K}$ must also hold for $Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ for any nonzero, nonunit $x$. In particular, if the set $\{[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\}$ is closed under the action of $G$, then we may apply our previous methods to $Cl_{K}/\left<[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\right$>. This is just another example of how misleading the quadratic case can be. As $\text{Gal}(K/\mathbb{Q})$ induces the subgroup $\{id,-id\}\leqslant\text{Aut}(Cl_{K})$ where $id$ is the identity and $id([I])=[I]^{-1}$, we see that any homomorphic image of $Cl_{K}$ can be realized by localizing at an integer—allowing significant restriction on $Cl_{K}$.
In general, it will be sufficient for a subgroup to be $\text{Aut}(Cl_{K})$-invariant. For example, if $G$ cannot admit a Galois action on $\mathbb{Z}/n\mathbb{Z}$, then we cannot have $Cl_{K}\cong\mathbb{Z}/an\mathbb{Z}$ for any $a\in\mathbb{N}$. This follows from the fact that $\mathbb{Z}/an\mathbb{Z}$ has a unique subgroup of order $n$. With this in hand, Theorem 3.4 becomes an immediate corollary of Theorem 3.1 in [7]. Furthermore, we can quickly improve upon Theorem 4.1.
Corollary 5.5.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, $Cl_{K}\not\cong\mathbb{Z}/np^{2}\mathbb{Z}$ for any $n\in\mathbb{N}$.
Once again, as we must have $h_{K}\equiv 0$ or $1$ mod $p$, this constitutes a serious restriction for extensions of (odd) prime degree. Now, note also that any Sylow $q$-subgroup of $Cl_{K}$ is invariant under automorphism. Thus, by the same logic, any restriction placed on $Cl_{K}$ may also be applied to its Sylow $q$-subgroups. Applying this also to Theorem 4.1, the following stronger result which appears in [12] is immediate.
Corollary 5.6.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, the Sylow $p$-subgroup of $Cl_{K}$ is not isomorphic to $\mathbb{Z}/p^{n}\mathbb{Z}$ for any $n\geqslant 2$.
In the same way, we can greatly strengthen Theorem 3.2.
Corollary 5.7.
Let $K$ be a Galois number field of degree $p^{r}$ and $S(q)$ a non-trivial Sylow $q$-subgroup of $Cl_{K}$. Then, $p=q$ or $|S(q)|\equiv 1$ mod $p$. Therefore, if $h_{k}=p^{r}q_{1}^{n_{1}}\dotsm q_{k}^{n_{k}}$, then $q_{i}^{n_{i}}\equiv 1$ mod $p$ for all $1\leqslant i\leqslant k$.
This was proven for abelian number fields by Fröhlich ([9]) which is a notably more restrictive condition. For example, observe that while Theorem 3.2 would not preclude an extension of degree $K:\mathbb{Q}]=3^{r}$ from having $h_{k}=55$, Corollary 5.7 shows this is impossible. In fact, such an extension cannot have $h_{K}=5n$ for any $n\in\mathbb{N}$ not divisible by $5$.
6. Overrings of Dedekind Domains
We conclude with a general consideration of overrings of Dedekind domains with eye toward developing a general criterion for when the class group of an overring of $\mathcal{O}_{K}$ is a Galois module via the action we have been studying. In particular, for an overring $\mathcal{O}_{K}\subseteq R\subseteq K$, we would like to know when the Galois group preserves fractional ideals of $R$ and thus admits the same norm-like group action on $Cl(R)$. We saw in Section 5 that even in the case of simple extensions, this may fail. Consider the following example.
Example 6.1.
Let $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-5}]$ and consider the overring $\mathcal{O}_{K}\left[\frac{1}{1+\sqrt{-5}}\right]$. Now, $\text{Gal}(K/\mathbb{Q})=\{\mathbf{1},\sigma\}$ where $\sigma$ represents complex conjugation. Note that $\sigma(R)$ fails to be an $R$-module as $\frac{1}{1+\sqrt{-5}}\cdot\sigma\left(\frac{1}{1+\sqrt{-5}}\right)=\frac{1}{1+\sqrt{-5}}\cdot\frac{1}{1-\sqrt{-5}}=\frac{1}{6}\notin\sigma(R)$.
The following is a classical result which characterizes all overrings of Dedekind domains with torsion class group. For details, the reader is encouraged to see [10].
Theorem 6.2.
Let $D$ be a Dedekind domain with torsion class group and $R$ an overring of $D$. Then
$R=\underset{\mathfrak{p}R\neq R}{\bigcap}D_{\mathfrak{p}}$
where the $\mathfrak{p}$ range over $\text{Spec}(D)$.
From this, we can quickly derive the following theorem of which we will make great use.
Theorem 6.3.
Let $D$ be a Dedekind domain with torsion class group. There exists a 111-1 correspondence between overrings of $D$ and nonempty subsets of $\text{Spec}(D)$ given by
$S\longleftrightarrow\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}.$
Proof.
By Theorem 6.2, every overring of $D$ can be expressed in the form $\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}$ for some $S\subseteq\text{Spec}(D)$. Thus, it suffices to show that $\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}\neq\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}$ for subsets $S_{1},S_{2}\subseteq\text{Spec}(D)$ whenever $S_{1}\neq S_{2}$. Without loss of generality, assume $\mathfrak{q}\in S_{1}\backslash S_{2}.$ Recall that $Cl(D)$ is torsion, so $|[\mathfrak{q}]|:=n$ is finite. Thus, $\mathfrak{q}^{n}=(q)$ for some $q\in D$. Now, $\mathfrak{q}$ is the unique prime ideal containing $q$, so $\frac{1}{q}\in\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}\backslash\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}$. ∎
Corollary 6.4.
Let $D$ be a Dedekind domain with torsion class group. Then $D\subseteq D[\frac{1}{a}]$ is a minimal ring extension if and only if $aD=\mathfrak{p}^{k}$ for some $k\in\mathbb{N}$.
Proof.
We prove the forward implication by contraposition. Assume there are distinct primes $\mathfrak{p},\mathfrak{q}\in\text{Spec}(D)$ which divide $aD$. Then, $a\in\mathfrak{p},\mathfrak{q}$, and hence $\mathfrak{p}D[\frac{1}{a}]=\mathfrak{q}D[\frac{1}{a}]=D$. Thus, by Corollary 6.3, there is an overring properly between $D$ and $D[\frac{1}{a}]$.
Now, for the reverse implication, assume $aD=\mathfrak{p}^{k}$ and $\mathfrak{q}D[\frac{1}{a}]=D[\frac{1}{a}]$ for some prime $\mathfrak{q}\in\operatorname{Spec}(D)$. Because $D=\mathfrak{p}^{k}, aa$ is strongly irreducible in $D$, and thus the saturation of $\{a^{n}\,|\,n\in\mathbb{N}\}$ in $D$ is simply $\{a^{n}\,|\,n\in\mathbb{N}\}$. Hence, $a^{k}\in\mathfrak{q}$ for some $k\in\mathbb{N}$ which implies $a\in\mathfrak{q}$. However, $\mathfrak{p}$ is the unique prime ideal containing $a$, so $\mathfrak{p}=\mathfrak{q}$. Therefore, $\mathfrak{p}$ is the unique prime ideal of $D$ such that $\mathfrak{p}D[\frac{1}{a}]=D[\frac{1}{a}]$. By the correspondence in Corollary 6.3, we have the $D\subseteq D[\frac{1}{a}]$ must be a minimal extension.
By Theorem 6.3, we know that any adjacent domain $D\subseteq D^{\prime}$, there is a unique prime $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}D^{\prime}=D^{\prime}$. What is more, any adjacent domain is a simple extension, and this $\mathfrak{p}$ determines the $\alpha$ such that $D^{\prime}=D[\alpha]$ as $\alpha=\frac{1}{a}$ where $aD=\mathfrak{p}^{n}$ where $|\mathfrak{p}|=n$. From these observations, we have the following Jordan-Hölder like result on overrings of Dedekind domains with finite class group.
Theorem 6.5.
Let $D$ be a Dedekind domain with finite group $Cl(D)$ such that every class contains a prime ideal. Let $R$ be an overring of $D$ which is minimal with respect to being a PID. If $D:=D_{0}\subseteq D_{1}\subseteq D_{2}\subseteq\dotsm\subseteq D_{r_{1}}:=R$ and $D:=D_{0}^{\prime}\subseteq D_{1}^{\prime}\subseteq D_{2}^{\prime}\subseteq\dotsm\subseteq D_{r_{2}}^{\prime}:=R$ are sequences of adjacent domains, then $r_{1}=r_{2}$ and the sequence of $\alpha_{i}\in R$ such that $D_{i+1}=D_{i}[\alpha_{i}]$ is unique up to reordering. What is more, such an overring is guaranteed to exists.
Proof.
We start by noting that for any adjacent domain $D\subseteq D_{1}$ there exists a unique $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}D_{1}=D_{1}$ by Theorem 6.3. Thus, we must have $D_{1}=\cap_{\mathfrak{q}\neq\mathfrak{p}}D_{\mathfrak{q}}=D[\frac{1}{a}]$ where $aD=\mathfrak{q}^{n}$ and $n=|[\mathfrak{q}]|$. Recall that $D[\frac{1}{a}]$ is the localization of $D$ at the complement of the set-theoretic union of all prime ideals of $D$ not containing $a$. In this case, that is all prime ideals omit $\mathfrak{q}$. By continuing this process and recalling the correspondence between prime ideals of a localization $D_{S}$ and prime ideals of $D$ disjoint from $S$ given by $\mathfrak{p}\mapsto\mathfrak{p}D_{S}$, we see that there is a unique, finite set of $r_{1}$ primes $\mathfrak{p}_{1},...,\mathfrak{p}_{r_{1}}\in\text{Spec}(D)$ such that $\mathfrak{p}R=R$. Thus, any sequence of adjacent domains between $D$ and $R$ must have length $r_{1}:=r$ and is unique up to our ordering of the $\mathfrak{p}_{i}$. In particular, if $\mathfrak{p}_{i}D_{i-1}$ is the unique prime such that $\mathfrak{p}_{i}D_{i}=D_{i}$, we must have $D_{i}=D_{i-1}[\frac{1}{a_{i}}]$ where $a_{i}D_{i-1}=\mathfrak{p}_{i}^{n_{i}}$.
Now, to construct such a PID and sequence, take some non-principal $\mathfrak{q}_{1}\in D$ such that $\mathfrak{q}_{1}^{n_{1}}=b_{1}D$, and let $D_{1}=D[\frac{1}{b}]$. As $\mathfrak{q}_{1}]\neq\text{Prin}(D)$, we must have $|Cl(D_{1})|<|Cl(D)|$. By the finiteness of the class group, if we continue this process, we will eventually have $|Cl(D_{t})|=1$ for some $t\in\mathbb{N}$ with $|Cl(D_{i})|\neq 1$ for all $0<i<t$. Thus, $D_{t}$ will be minimal with respect ot being a PID. ∎
This theorem tells us that any ring of integers in particular is finitely many adjacent steps away from being a PID.
Example 6.6.
Recall $Cl_{K}\cong\mathbb{Z}/4\mathbb{Z}$ for $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-14}]$. Via the process described in the proof of Theorem 6.5, $\mathbb{Z}[\sqrt{-14}]\subseteq\mathbb{Z}[\sqrt{-14},\frac{1}{2}]\subseteq\mathbb{Z}[\sqrt{-14},\frac{1}{6}]$ is a series of adjacent domains with $\mathbb{Z}[\sqrt{-14},\frac{1}{6}]$ a minimal PID over $\mathbb{Z}[\sqrt{-14}]$.
Now, in the case of rings of integers, $R$ being a minimal PID over $\mathcal{O}_{K}$ does not imply the existence of such chain of adjacent domains. In fact, the following example implies that there always exists a PID over $\mathcal{O}_{K}$ failing to have this property except in the case that the ring of integers itself is a PID.
Example 6.7.
Let $\mathcal{O}_{K}$ be a ring of integers with $Cl_{K}=\mathbb{Z}/2\mathbb{Z}$, and let $P$ be the set of all non-principal prime ideals of $\mathcal{O}_{K}$. Now, taking $S=\cup_{\mathfrak{p}\in P}\mathfrak{p}$, we have that $R_{S^{C}}=\{\frac{r}{a}\,|\,r\in R,a\notin S\}$ also has class group $\mathbb{Z}/2\mathbb{Z}$ by Theorem 1.5 of [3]. Now, take a (non-principal) prime $\mathfrak{q}\in\text{Spec}(R)$, and let $bR_{S}=\mathfrak{q}^{n}$ for some nn\ $\in\mathbb{N}$. Then, $R_{S}[\frac{1}{b}]$ is minimal over $\mathcal{O}_{K}$ with respect to being a PID, but there is no finite chain of adjacent domains between them by Theorem 6.3.
We now return to the particular case of rings of integers. For a more general treatment of overrings of Dedekind domains, the reader is referred to [3] and [4]. In the following theorem, we adopt the notation of [15] and denote the group of fractional ideals of a domain $R$ as $G(R)$.
Theorem 6.8.
Let $K$ be a Galois number field with $G:=\text{Gal}(K/\mathbb{Q})$ and $R$ an overring of the ring of integers $\mathcal{O}_{K}$. The following are equivalent:
  1. (1)
    The action $\sigma\cdot[I]=[\sigma(I)]$ for $\sigma\in G$ and $I]\in G(R)$ is well-defined.
  2. (2)
    $R$ is Galois-invariant. That is for all $\sigma\in G$, $\sigma(R)=R$.
  3. (3)
    For all $\mathfrak{p}\in\text{Spec}(\mathcal{O}_{K})$ and $\sigma\in G$, $\mathfrak{p}R=R\implies\sigma(\mathfrak{p})R=R$.
  4. (4)
    For any $\mathfrak{p},\mathfrak{q}\in\text{Spec}(\mathcal{O}_{K})$ lying over the same rational prime $p$, $\mathfrak{p}R=R\implies\mathfrak{q}R=R$.
Proof.
$(1\Rightarrow 2)$ By well-definedness, we must have $\sigma(R)$ an $R$-module for all $\sigma\in G$. Then, for all $r\in R$, $r\cdot\sigma(1)=r\cdot 1=r\in\sigma(R)$, so we have $R\subseteq\sigma(R)$. By the same argument, $R\subseteq\sigma^{-1}(R)$ and thus $\sigma(R)\subseteq R$. Therefore, $\sigma(G)=G$ as desired.
$(2\Rightarrow 1)$ Assume $R$ is Galois-invariant. Let $I\in G(R)$ and $\sigma\in G$. By definition, we have $\alpha I\subseteq R$ for some $\alpha\in K$. By our assumption, $\sigma(\alpha)\sigma(I)=\sigma(\alpha I)\subseteq\sigma(R)\subseteq R$ for $\sigma(\alpha)\in K$. Now, let $r\in R$, then $r^{\prime}:=\sigma^{-1}(r)\in R$ by assumption. Therefore, $r\sigma(I)=\sigma(r^{\prime})\sigma(I)\in\sigma(r^{\prime}I)\subseteq\sigma(I)$ as $I$ is an $R$-module. We conclude that $\sigma(I)$ is a fractional ideal, and thus the map $I\mapsto\sigma(I)$ from $G(\mathcal{O}_{K})$ to $G(R)$ descends to the group action in $1$.
$(2\Rightarrow 3)$ We will proceed by contraposition. Assume that there exists $\mathfrak{p}\in\text{Spec}(\mathcal{O})_{K}$ and $\sigma\in G$ such that $\mathfrak{p}R=R$ but $\sigma(\mathfrak{p})R\neq R$. Then there exists $\frac{a}{b}\in U(R)\cap\mathfrak{p}$, so $\frac{b}{a}\in R$. However, $\sigma(\frac{b}{a})=\frac{\sigma(b)}{\sigma(a)}\notin R$ because $\sigma(\frac{a}{b})=\frac{\sigma(a)}{\sigma(b)}\in\mathfrak{q}$ and $\mathfrak{q}\cap U(R)=\emptyset$. Therefore, $R$ is not Galois-invariant.
$(3\Rightarrow 2)$ We will once again proceed by contraposition. Assume there exists $\frac{a}{b}\in R$ with $a,b\in\mathcal{O}_{K}$ and $\sigma\in G$ such that $\sigma(\frac{a}{b})=\frac{\sigma(a)}{\sigma(b)}\notin R$. Recall that $R=\cap_{\mathfrak{p}R\neq R}(\mathcal{O}_{K})_{\mathfrak{p}}$. Then, there exists some $\mathfrak{q}\in\text{Spec}(R)$ such that $\mathfrak{q}R\neq R$ with $\sigma(b)\in\mathfrak{q}$. Now, $b=\sigma^{-1}(\sigma(b))\in\sigma^{-1}(\mathfrak{q}):=\mathfrak{q}^{\prime}$ implies $\mathfrak{q}^{\prime}R=R$. Thus, we have $\mathfrak{q}^{\prime}\in\text{Spec}(R)$ such that $\mathfrak{q}^{\prime}R=R$, but $\sigma(\mathfrak{q}^{\prime})R=\mathfrak{q}R\neq R$.
$(3\Leftrightarrow 4)$ The equivalence of these conditions follows directly from the transitive action of $G$ on the primes lying over a given rational prime (for details, see [13]). This completes the proof. ∎
This theorem demonstrates the fact that any class group of an overring of $\mathcal{O}_{K}$ which admits such a group action can be realized an a simple extension of the form $\mathcal{O}_{K}[\frac{1}{n}]$ where $n$ is an integer. In particular, we see that the types of extensions studied in the previous section are sufficient to achieve all homomorphic images which are of use to us in restricting the structure of $Cl_{K}$.