### v4

# Galois Action and Localization in Number Fields

Abstract.
For a Galois number field $K$, the Galois group $\text{Gal}(K/\mathbb{Q})$ acts on the class group $Cl_{K}$ in a very natural way: $\sigma\cdot[I]=[\sigma(I)]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$, $I]\in Cl_{K}$. In this paper, we will explore how the unique properties of this group action work together to elucidate the relationship between these two groups—developing and expanding upon some known results from a new perspective. To this end, we explore the class groups of localizations of the ring of integers $\mathcal{O}_{K}$. These turn out to be powerful tools for understanding $Cl_{K}$ and overrings of $\mathcal{O}_{K}$. The paper concludes with some interesting observations about normset arithmetic—a topic intimately related to this action.
1. Introduction and Notation
Throughout this paper $K$ will denote a Galois number field with Galois group $G:=\text{Gal}(K/\mathbb{Q})$, and $\mathcal{O}_{K}$ its ring of integers with class group $Cl_{K}$ and class number $h_{K}=|Cl_{K}|$. The Galois group of $K$ acts on the class group in a very natural way. For any $\sigma\in G$ and $I]\in Cl_{K}$, we can define the action $\sigma\cdot[I]=[\sigma(I)]$. This is a well-defined group action which provides us with some tools for characterization the relationship between the two groups. More than this, we also have the peculiar property that $\sigma\cdot([I][J])=(\sigma\cdot[I])(\sigma\cdot[J])$ for any $\sigma\in G$ and $I],[J]\in Cl_{K}$. This induces a map from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\mapsto\bar{\sigma}$ where $\bar{\sigma}([I])=[\sigma(I)]$. Finally, we have the “norm property” that $\prod_{\sigma\in G}\sigma\cdot[I]=\left[\prod_{\sigma\in G}\sigma(I)\right]=\text{Prin}(\mathcal{O}_{K})$ for any $I]\in Cl_{K}$—for details, see [22]. Note, in this paper we will study the case where $K$ is Galois over $\mathbb{Q}$ with the knowledge that the theorems herein apply also to the case when $K/L$ is a Galois extension of number fields with $\mathcal{O}_{L}$ a principal ideal domain (PID). Now, these characteristics of the action motivate the following definition.
Definition 1.1.
Let $G$ and $A$ be groups with $A$ abelian, and let
$\alpha:G\times A\longrightarrow A,\quad(g,a)\mapsto g\cdot a$
be a map. If $\alpha$ satisfies the following properties:
  1. (1)
    $g_{1}\cdot(g_{2}\cdot a)=(g_{1}g_{2})\cdot a$ for all $g_{1},g_{2}\in G$, $a\in A$,
  2. (2)
    $e_{G}\cdot a=a$ for all $a\in A$,
  3. (3)
    $g\cdot(a_{1}a_{2})=(g\cdot a_{1})(g\cdot a_{2})$ for all $g\in G$, $a_{1},a_{2}\in A$,
  4. (4)
    $\prod_{g\in G}(g\cdot a)=e_{A}$ for all $a\in A$,
then we say that $\alpha$ is a norm-like action.
Throughout this paper, we will see how properties 1-4 can be used to place strong restrictions on the structure of $Cl_{K}$ given $G$ and vice versa. We will place a particular emphasis on techniques developed and some applications to factorization theory. While the class group is inherently connected to the arithmetic of rings of integers, this particular action is intimately related to the factorization properties of the normset of a ring of integers. This connection was first studied by the first author in [11]. The recent work of Geroldinger, Halter-Koch, and Zhong in [15] has inspired renewed interest in this action as it relates to normset factorization—particularly in the quadratic case.
Throughout this paper, we will present new proofs of some existing theorems as well as extend various results on the structure of $Cl_{K}$ using new techniques. Previous work done on this subject has been primarily through the lens of representation theory—considering $Cl_{K}$ as a $G$-module. This approach was utilized by Fröhlich in 1952 ([14]) and subsequently many others. Cornell and Rosen used this approach in [10] to give results on the structure of $Cl_{K}$ when $h_{K}$ is known. Others such as Lemmermeyer ([21]) and Iwasawa ([19]) used the $G$-module structure of $Cl_{K}$ along with assumptions on intermediate fields to give various results on the $p$-rank of the class group. In [20], Komatsu and Nakano generalize this work, using the Galois-module structure of the class group and some group theoretic considerations, to give further restrictions on the $p$-rank of $Cl_{K}$—once again using assumptions the class group of intermediate fields. We will develop some of the tools used in this paper from a strictly group-theoretic point of view. However, we diverge from these methods by considering the class groups of overring of $\mathcal{O}_{K}$ rather than that of intermediate rings. For a thorough review of the previous work done in this area, see [23].
In the following section, we explore how conditions 1 and 2, the typical conditions of a group action, together with 4 place restrictions of the structure of $Cl_{K}$ for a Galois number field $K$ of given degree. In Section 3, we focus on applications of condition 3 to the inverse class group problem. In particular, which number fields $K$ can have a given abelian group as their class group. Section 4 considers the structure of the class group of localizations of the form $\mathcal{O}_{K}[\frac{1}{\alpha}]$. This ultimately strengthens many of the previously developed results by allowing us to restrict to Sylow $p$-subgroups of the class group. In Section 5, the authors consider the class groups of overrings of Dedekind domains broadly. The paper culminates with a characterization of those overrings of Galois rings of integers whose class groups admit the same norm-like group action described above. This ultimately demonstrates that all class groups of interest can be realized as the class group of a simple localizations. Finally, Section 6 considers the arithmetic of norms of algebraic rings of integers—a topic closely tied to this action. We conclude by drawing an interesting connection between this arithmetic and the partition problem.
2. Direct Consequences of the Group Action
In this section, we produce results on the structure of the class group of Galois number fields which follow directly from the action of the Galois group $G=\text{Gal}(F/\mathbb{Q})$ on the class group $Cl_{K}$ in conjunction with the norm property $\prod_{\sigma\in G}\sigma(I)\in\text{Prin}(\mathcal{O}_{K})$ for any ideal $I\subseteq\mathcal{O}_{K}$. We begin with a lemma which we will use extensively throughout this paper.
Lemma 2.1.
Let $K$ be a Galois number field with $G=\text{Gal}(K/\mathbb{Q})$ and class group $Cl_{K}$. If $G$ acts trivially on $I]\in Cl_{K}$, the order of $[I]$ divides $|G|=[K:\mathbb{Q}]$.
Proof.
Assume $G$ acts trivially on $I]\in Cl_{K}$. Now, by the norm property, we have
$\prod_{\sigma\in G}\sigma(I)\in\text{Prin}(\mathcal{O}_{K}).$
Thus, we have
$\prod_{\sigma\in G}[\sigma(I)]=\text{Prin}(\mathcal{O}_{K})\Rightarrow{}\prod_{\sigma\in G}\sigma\cdot[I]=\text{Prin}(\mathcal{O}_{K}).$
As $G$ acts trivially on $[I]$, we have
$\text{Prin}(\mathcal{O}_{K})=\prod_{\sigma\in G}\sigma\cdot[I]=\prod_{\sigma\in G}[I]=[I]^{|G|}.$
This completes the proof. ∎
The first result we present follows from Lemma 2 in [10] using the representation theoretic approach. We will prove it directly here and generalize the result in section $4$ with the techniques developed therein.
Theorem 2.2.
Let $K$ be a Galois number field of degree $p^{r}$ where $p$ is prime. Then, $h_{K}\equiv 0\text{ or }1\text{ mod }p$.
Proof.
Assume that $h_{K}\not\equiv 1$ mod $p$. Once again, let $G$ act on $Cl_{K}$ via the action $\sigma\cdot[I]=[\sigma(I)]$. By the orbit-stabilizer theorem, the lengths of the orbits of this group action divide $|G|=[K:\mathbb{Q}]=p^{r}$. In particular, they must be elements of the set $\{1,p,p^{2},\dots,p^{r}\}$.
Now, the identity of $Cl_{K}$ has a trivial orbit of order $1$. The orbits partition $Cl_{K}$, and all non-trivial orbits have order congruent to 0 mod $p$. Hence, as we assumed $h_{K}\not\equiv 1$ mod $p$, there must be some non-identity class of $Cl_{K}$ in a trivial orbit. Call this element $J]\neq\text{Prin}(\mathcal{O}_{K})$. By Lemma 2.1, $[J]$ has order dividing $p^{r}$, and as we assumed $[J]$ was not the identity, $p$ must divide the order of $[J]$. Therefore, by Lagrange’s theorem, $p\text{ divides }h_{K}$, so $h_{K}\equiv 0$ mod $p$. This completes the proof.
This theorem demonstrates one of the many ways in which quadratic number fields are exceptional. Beyond their peculiarity of being Galois in general, Theorem 2.2 leaves all class numbers permissible in the quadratic case. The same is true of the following.
Theorem 2.3.
If $K$ is a Galois number field of degree $n$ and $p$ the smallest prime divisor of $n$, $h_{k}=1$ or $h_{k}\geqslant p$.
Proof.
Assume for the purpose of contradiction that $1<h_{k}<p$. As $h_{k}<p$, the smallest prime divisor of $|G|$, the orbit-stabilizer theorem tells us every element of $Cl_{K}$ has a trivial orbit under the action of $G$ on $Cl_{K}$. As $h_{k}>1$, we may take $J]\neq\text{Prin}(\mathcal{O}_{K})$ in $Cl_{K}$. Then, by Lemma 2.1, we must have $J]^{|G|}=\text{Prin}(\mathcal{O}_{K})$, so the order of $[J]$ divides $|G|=n$. However, this is a contradiction as $1<|[J]|<h_{k}<p$, and $p$ is the smallest prime divisor of $n$. ∎
The question of whether every finite abelian group can be realized as the class group of a quadratic ring of integers is a famous open problem. We will consider some related questions in the next section. For now, we now turn our attention to some results on Galois number fields of odd degree.
Theorem 2.4.
Let $K$ be a Galois number field of odd degree. Then, $Cl_{K}\ncong\mathbb{Z}/2n\mathbb{Z}$ for any $n\in\mathbb{N}$.
Proof.
Assume for the purpose of contradiction that $Cl_{K}\cong\mathbb{Z}/2n\mathbb{Z}$. Then, $Cl_{K}$ has a unique element of order 2, call it $[J]$. Thus, as each $\sigma\in G$ is an automorphism, we must have $\sigma\cdot[J]=[\sigma(J)]=[J]$ for all $\sigma\in G$. Hence, $G$ acts trivially on $[J]$, so by Lemma 2.1 the order of $[J]$ divides $|G|$. However, this is a contradiction as $[J]$ has order $2$ and $|G|$ is odd. ∎
In particular, for a Galois number field of odd degree, $Cl_{K}$ cannot have a unique invariant factor of even order. Notably, this precludes cyclic groups of even order. Now, from Theorem 2.4 we get the following factorization result which first appears as Lemma 3.5 in [12].
Corollary 2.5.
Let $K$ be a Galois number field of odd degree. Then $\mathcal{O}_{K}$ is an HFD if and only if $\mathcal{O}_{K}$ is a UFD.
Proof.
A well-known result from Carlitz ([6]) tells us $\mathcal{O}_{K}$ is an HFD if and only if $h_{K}=1$ or $2$. Noting that Theorem 2.4 disallows $h_{K}=2$ and $\mathcal{O}_{K}$ is a UFD if and only if $h_{K}=1$ completes the proof. ∎
Note that this result also follows directly from Theorem 2.3. We conclude this section with a specific application of these results to factorization in cyclotomic number fields. These number fields in particular played a central role in the development of modern factorization theory. In his attempt to prove Fermat’s Last Theorem, Ernst Kummer erroneously assumed that the cyclotomic ring of integers $\mathbb{Z}[\xi_{p}]$ was a unique factorization domain for any primitive $p^{th}$ root of unity. It was Dedekind’s remarkable observation that while unique factorization for elements was lost, unique factorization of ideals was retained for all rings of integers. Thus, we are able to turn questions about the factorization of elements in rings of integers into questions about ideals and, in turn, the class group. Being Galois in general, the cyclotomic number fields provide an apt setting for the applications of the results we have developed in this section.
Theorem 2.6.
Let $p<23$ be an odd prime and $a\in\mathbb{Z}$. If $K$ is the splitting field of $x^{p}-a$, $\mathcal{O}_{K}$ is an HFD if and only if $\mathcal{O}_{K}$ is a UFD.
Proof.
We begin by noting that $a=1$ is a degenerative case in which the theorem holds trivially. Now, as the splitting field of $x^{p}-a$, $K$ is Galois over $\mathbb{Q}$. Furthermore, we have $K\cong\mathbb{Q}(\omega,\alpha)$ where $\omega$ is a primitive $p^{th}$ root of unity and $\alpha$ a root of $x^{p}-a$. This gives us the following lattice.
$\mathbb{Q}$$\mathbb{Q}(\omega)$$\mathbb{Q}(\alpha)$$\mathbb{Q}(\omega,\alpha)$
Where $\mathbb{Q}(\omega,\alpha)$ is Galois over $\mathbb{Q}(\omega)$ of degree $p$. Now, as $p<23$, $\mathbb{Z}[\omega]$ is a UFD, so the result follows from Corollary 2.5.
[…]
Recall that the action of $G=\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ induces a homomorphism from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\overset{\psi}{\mapsto}\phi$ where $\phi([I])=\sigma\cdot[I]=[\sigma(I)]$. Note, the kernel of this homomorphism is precisely the kernel of the action. In this section, we will see how this map can be used to further restrict the structure of $Cl_{K}$ for a Galois number field $K$. Conversely, we will use this tool to help answer questions related to the inverse class group problem. This terminology is used in the spirit of the inverse Galois problem. In particular, the inverse class group problem asks if an arbitrary abelian group can be realized as the class group of a ring of integers. This question was famously answered in the positive for the more general class of Dedekind domains by Claborn in [9]. In the case of imaginary quadratics, the answer is known to be no. Certain groups can be shown by brute force not to appear, the smallest among them being $(\mathbb{Z}/3\mathbb{Z})^{3}$. For details, the reader is encouraged to see [25]. The question remains open for rings of integers in general. In this section, for a given finite abelian group $A$, we will explore which rings of integers may admit $Cl_{K}\cong A$. We begin with a direct proof of another result from [10] which we will return to with some new found tools in section 4.
Theorem 3.1.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, $Cl_{K}\ncong\mathbb{Z}/p^{n}\mathbb{Z}$ for $n\geqslant 2$.
Proof.
[…]
Assume $H$ is trivial. Then, each element of $G$ acts trivially on the class group. Therefore, by Lemma 2.1, every element of $Cl_{K}$ has order dividing $|G|=p$, but this contradicts our assumption that $Cl_{K}\cong\mathbb{Z}/p^{n}\mathbb{Z}$ with $n\geqslant 2$. Hence, $H$ is the unique subgroup of order $p$.
[…]
$\prod_{\sigma\in G}\sigma(J)\in\text{Prin}(\mathcal{O}_{K})$
implies
$\prod_{\sigma\in G}[\sigma(J)]=\prod_{\sigma\in G}\sigma[(J)]=\prod_{i=0}^{p-1}\psi^{i}([J])=\text{Prin}(\mathcal{O}_{K})$
Furthermore, $\psi^{i}([J])=(\gamma^{(p-1)p^{n-2}})^{i}([J])=[J]^{\left(\alpha^{(p-1)p^{n-2}}\right)^{i}}$. Note, $(p-1)p^{n-2}=\varphi(p^{n-1})$, so $\alpha^{(p-1)p^{n-2}}\equiv 1$ mod $p^{n-1}$. Also, because $\alpha$ is a primitive root mod $p^{n}$, each $(\alpha^{(p-1)p^{n-2}})^{i}=\alpha^{i(p-1)p^{n-2}}$ is unique mod $p^{n}$ for each $0\leqslant i\leqslant p-1$. Hence, they are precisely the elements $\{1,p^{n-1}+1,2p^{n-1}+1,\dots,(p-1)p^{n-1}+1\}$ mod $p^{n}$. Taking their sum, we get
$\sum_{i=0}^{p-1}\left(1+i\cdot p^{n-1}\right)=p+p^{n-1}\sum_{i=1}^{p-1}i=p+p^{n-1}\cdot\frac{(p-1)p}{2}=p+\frac{p-1}{2}\cdot p^{n}\equiv p\text{ mod }p^{n}$
Hence,
$\prod_{i=0}^{p-1}\psi^{i}([J])=\prod_{i=0}^{p-1}[J]^{\left(\alpha^{(p-1)p^{n-2}}\right)^{i}}=[J]^{\sum_{i=0}^{p-1}1+i\cdot p^{n-1}}=[J]^{p}=\text{Prin}(\mathcal{O}_{K})$
But this implies the order of $[J]$ divides $p$, and $[J]$ is of order $p^{n}$ with $n\geqslant 2$. This is a contradiction. Therefore, $Cl_{K}$ cannot be $\mathbb{Z}/p^{n}\mathbb{Z}$ for any $n\geqslant 2$.
Recall that Theorem 2.2 tells us (in particular) that for a Galois number field of degree $p$, we must have $h_{k}\equiv 0$ or $1$ mod $p$. Thus, Theorem 3.1 gives us a further restriction in the first case on the structure of the class group. We now shift our attention to the inverse class group. First, we consider the case when $h_{k}$ is assumed to be an odd prime rather than $K:\mathbb{Q}]$. Note, this is equivalent to asking which number fields $K$ can have $Cl_{K}\cong\mathbb{Z}/p\mathbb{Z}$.
Theorem 3.2.
Let $K$ a Galois number field with $n=[K:\mathbb{Q}]$ and $h_{K}=p$ prime. Then, $p|n$ or gcd(p1,n)>1(p-1,n)>1.
Proof.
[…]
This theorem tells us that having class group $\mathbb{Z}/p\mathbb{Z}$ is a relatively restrictive condition. For example, only Galois number fields of degree divisible by $2$ or $17$ can have class group $\mathbb{Z}/17\mathbb{Z}$. We continue this section with a few specific examples of the inverse class group problem which present new techniques and highlight the interesting interplay between the induced homomorphism and the norm property of the group action.
Example 3.3.
Let us assume $Cl_{K}\cong\mathbb{Z}/13\mathbb{Z}$ for some Galois number field $K$. None of the theorems developed thus far preclude a cubic Galois extension from having such a class group, and indeed we see that the number field with defining polynomial $x^{3}-x^{2}-354x-2441$ is one such example. Now, any cubic Galois extension has $G\cong\mathbb{Z}/3\mathbb{Z}$, so the induced homomorphism $\psi:G\rightarrow\text{Aut}(Cl_{K})$ must be injective as the alternative would imply there exists an element in $\mathbb{Z}/13\mathbb{Z}$ of order $|G|=3$.
[…]
The following theorem is inspired by this example and demonstrates that the phenomenon described at the end must occur in general for $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$.
Theorem 3.4.
Let $p$ be prime and $\phi=(a_{1},a_{2},...,a_{p-1})$ a cyclic generator of $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$. If $n$ properly divides p1p-1, then the elements in each disjoint cycle of $\phi^{n}$ sum to 0 mod $p$.
Proof.
Say $\phi(1)=a$, then we must have $\phi(x)=a\cdot x$ for any $x\in\mathbb{Z}/p\mathbb{Z}$. As $\phi$ generates $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$, this also implies that $a$ is a primitive root mod $p$. Now, any $x\in\mathbb{Z}/p\mathbb{Z}$ is contained in a (p1)/n(p-1)/n cycle in $\phi^{n}$. Thus, the cycle is of the form $(x\,\,\,xa^{n}\,\,\,xa^{2n}\dotsm xa^{(\frac{p-1}{n}-1)n})$. Taking the sum, we get
$x\sum_{r=0}^{\frac{p-1}{n}-1}(a^{n})^{r}=x\cdot\frac{1-(a^{n})^{\frac{p-1}{n}}}{1-a^{n}}=x\cdot\frac{1-a^{p-1}}{1-a^{n}}$
By Fermat’s Little Theorem, we have $1-a^{p-1}\equiv 0$ mod $p$. The result then follows from the fact that $n$ properly divides p1p-1, so $1-a^{n}\not\equiv 0$ mod $p$.
Example 3.5.
Consider a Galois number field $K$ with class group $Cl_{K}\cong(\mathbb{Z}/2\mathbb{Z})^{3}$. It is well known that $\text{Aut}((\mathbb{Z}/2\mathbb{Z})^{3})\cong\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$, the group of invertible $3\times 3$ matrices over the field of two elements with order $|\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})|=168=2^{3}\cdot 3\cdot 7$. Employing the same methods as in the proof of Theorem 3.2, we immediately see that the order of the extension $n=[K:\mathbb{Q}]$ must be divisible by $2,3$ or $7$. However, Theorem 2.2 tells us that no Galois extension of degree $3^{r}$ can have class group $(\mathbb{Z}/2\mathbb{Z})^{3}$ for any $r\geqslant 1$. We will now show that in fact no Galois number field of degree $3m$ where $\text{gcd}(2,m)=\text{gcd}(7,m)=1$ can have class group $(\mathbb{Z}/2\mathbb{Z})^{3}$.
Assume $G=3m$ where $m$ is not divisible by $2$ or $7$. Once again, the homomorphism $\psi:G\rightarrow\text{Aut}(Cl_{K})\cong\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ must be non-trivial as $Cl_{K}$ has no non-identity elements of order dividing $|G|=3m$. Thus, $im(\psi)$ must be a subgroup of order $3$ by the first isomorphism theorem. Now, consider the following element of order $3$ in $\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ and its powers:
$E=\begin{bmatrix}1&0&0\\ 0&0&1\\ 0&1&1\end{bmatrix}\hskip 14.22636ptE^{2}=\begin{bmatrix}1&0&0\\ 0&1&1\\ 0&1&0\end{bmatrix}\hskip 14.22636ptE^{3}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
Now, $E$ has characteristic polynomial $p_{E}(x)=(x+1)(x^{2}+x+1)$, so we see it has eigenvalue $\lambda=1$. From here, one can show $(1,0,0)^{T}$ is a corresponding eigenvector. That is, $E$ fixes the vector $(1,0,0)^{T}$, and so $E^{2}$ must also. Furthermore, by the second Sylow theorem and the fact that similar matrices have the same eigenspaces, we conclude that any element of order $3$ in $\text{GL}_{3}(\mathbb{Z}/2\mathbb{Z})$ fixes $(1,0,0)^{T}$. Therefore, as $im(\psi)$ is a subgroup of order $3$, we must have $(1,0,0)\in(\mathbb{Z}/2\mathbb{Z})^{3}$ fixed by every element of $G$, but this implies the order of $(1,0,0)$, namely $2$, divides $|G|=3m$, a contradiction. Therefore, we see that some degrees such as $K:\mathbb{Q}]=15$ which were not previously disallowed are impossible for $Cl_{K}\cong(\mathbb{Z}/2\mathbb{Z})^{3}$.
Similar methods can be employed to investigate the inverse class group problem for $p$-elementary abelian groups as well as cyclic groups of order $2^{m}p^{r}$ with $m\in\{0,1\}$ and $r\geqslant 1$ where $p$ is an odd prime.
4. Localization and the Class Group
In this section, we will use localization to strengthen some of the results developed previously. First, recall that any overring—in particular, any localization—of a Dedekind domain remains Dedekind. Intuitively, as a localization of $R$ is attained by turning a set of elements $S\subseteq R$ into units, we expect the class group of $R_{S}$ to be smaller than that of $R$ as those non-principal ideals $I\subseteq R$ for which $I\cap S\neq\emptyset$ will become principal. As we will demonstrate, “small” localizations will tend to lead to small changes in the class group. In the context of rings of algebraic integers, this will be useful in further restricting the possible structure of $Cl_{K}$. The following result can be viewed as a particular case of Theorem 1.5 in [8]—one of Claborn’s pioneering works on Dedekind domains.
Theorem 4.1.
Let $D$ be a Dedekind domain and $\alpha\in D$ a nonzero nonunit with
$(\alpha)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}.$
Then $Cl(D[\frac{1}{\alpha}])\cong Cl(D)/\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$>.
Note, $D[\frac{1}{\alpha}]$ is the localization of $D$ at the multiplicative set $\{\alpha^{n}\,|\,n\in\mathbb{N}\}$ which is the complement of the set-theoretic union of all prime ideals not containing $\alpha$.
Proof.
Let $\psi:Cl(D)\rightarrow Cl(D[\frac{1}{\alpha}])$ be the canonical map $I]\overset{\psi}{\mapsto}\left[ID[\frac{1}{\alpha}]\right]$. By Propositions 1.3 and 1.5 of [8], $\psi$ maps $Cl(D)$ onto $Cl(D[\frac{1}{\alpha}])$ and the kernel is generated by the classes $\mathfrak{p}_{\beta}]$ where $\mathfrak{p}_{\beta}$ ranges over the prime ideas such that $\mathfrak{p}_{\beta}D[\frac{1}{\alpha}]=D[\frac{1}{\alpha}]$. As $D[\frac{1}{\alpha}]=D_{S}$ where $S=\{\alpha^{n}\,|\,n\in\mathbb{N}\}$, these are precisely the prime ideals $\mathfrak{p}$ such that $\mathfrak{p}\cap S\neq\emptyset$ or the primes containing $\alpha^{n}$ for some $n\in\mathbb{N}$. By observing that these are the prime divisors of $(\alpha)$, we obtain the result by the first isomorphism theorem.
Notably, given $Cl(D)$ is torsion, Theorem 4.1 implies that any homomorphic image of $Cl(D)$ can be realized as the class group of such a localization of $D$. Claborn gives a similar, slightly weaker result in [8] where, given a Dedekind domain $A$ with arbitrary class group $G$ and subgroup $H$, he constructs a Dedekind polynomial ring over $A$ with class group isomorphic to $G/H$. As he answered the inverse class group problem for Dedekind domains, Theorem 4.1 brings us a step closer for overrings of $\mathcal{O}_{K}$. In essence, $\mathcal{O}_{K}[\frac{1}{\alpha}]$ is very nearly a ring of integers itself. In fact, it is the precise analog if the base ring $\mathbb{Z}$ were replaced by $\mathbb{Z}[\frac{1}{N(\alpha)}]$ for some nonzero $\alpha\in\mathcal{O}_{K}$.
Theorem 4.2.
Let $K$ be a Galois number field and $\alpha\in\mathcal{O}_{K}$ a nonzero element. Then $\mathcal{O}_{K}[\frac{1}{\alpha}]$ is integral over $\mathbb{Z}[\frac{1}{N(\alpha)}]$.
Proof.
Note, it is sufficient to show $\frac{1}{\alpha}$ is integral over $\mathbb{Z}[\frac{1}{N(\alpha)}]$. Let $f(x)\in\mathbb{Z}[x]$ be the minimal polynomial of $\alpha$. Noting that $N(\alpha)$ is the product of the roots of $f(x)$, we write $f(x)=x^{n}+m_{n-1}x^{n-1}+\dotsm+m_{1}x+N(\alpha)$. Thus,
$N(\alpha)+m_{1}\alpha+\dotsm+m_{n-1}\alpha^{n-1}+\alpha^{n}=0.$
Multiplying by $\frac{1}{N(\alpha)\alpha^{n}}$ yields
$\left(\frac{1}{\alpha}\right)^{n}+\frac{m_{1}}{N(\alpha)}\left(\frac{1}{\alpha}\right)^{n-1}+\dotsm+\frac{m_{n-1}}{N(\alpha)}\left(\frac{1}{\alpha}\right)+\frac{1}{N(\alpha)}=0.$
This completes the proof. ∎
From this, we can see that $\mathcal{O}_{K}[\frac{1}{\alpha}]$ is in fact the integral closure of $\mathbb{Z}[\frac{1}{N(\alpha)}]$ in $K$ as the former is integrally closed.
$\mathbb{Q}$}K{K}[$\mathbb{Z}\!\left[\frac{1}{N(\alpha)}\right]$$\mathcal{O}_{K}\!\left[\frac{1}{\alpha}\right]$}
Note also that $Cl(\mathcal{O}_{K}[\frac{1}{\alpha}])$ is finite and retains the useful property that each class contains a prime ideal. This follows directly from the correspondence between prime ideals of $\mathcal{O}_{K}[\frac{1}{\alpha}]$ and prime ideals of $\mathcal{O}_{K}$ disjoint from $\{\alpha^{n},\,n\in\mathbb{N}\}$. This allows for the application of many theorems on rings of integers to be applied to $\mathcal{O}_{K}[\frac{1}{\alpha}]$. The following gives one such example.
Example 4.3.
Consider the ring of integers $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-14}]$ with class group $Cl_{K}\cong\mathbb{Z}/4\mathbb{Z}$ generated by $\mathfrak{p}=(3,1-\sqrt{-14})$. Writing $\mathfrak{q}=(2,\sqrt{-14})$, it is not difficult to show $\mathfrak{q}]=[\mathfrak{p}^{2}]$. Now $\mathfrak{q}^{2}=(2)$, so by Theorem 4.1 we have $Cl(\mathcal{O}_{K}[\frac{1}{2}])\cong\mathbb{Z}/2\mathbb{Z}$ implying $\mathcal{O}_{K}[\frac{1}{2}]=\mathbb{Z}[\sqrt{-14},\frac{1}{2}]$ is an HFD which is not a UFD.
Most significantly, Theorem 4.1 allows us to excise the minimal amount of $Cl_{K}$ and thus produce relatively large homomorphic images. In fact, we can realize any homomorphic image of $Cl_{K}$ in this way. In general, if $\gamma:Cl_{K}\rightarrow H$ is a homomorphism with $\text{ker}(\gamma)=\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{n}]\right$>, $Cl(\mathcal{O}_{K}[\frac{1}{\alpha}])\cong H$ where $(\alpha)=(\mathfrak{p}_{1}\mathfrak{p}_{2}\dotsm\mathfrak{p}_{n})^{h_{K}}$.
The remainder of this section will be dedicated to leveraging this fact to extend the techniques and theorems developed in sections 2 and 3. Our hope is that the Galois group of $K$ will act on the class group of $\mathcal{O}_{K}[\frac{1}{\alpha}]$ in the same manner as $Cl_{K}$ which would allow us to place the same restrictions on the homomorphic images of the class group—further constraining the structure of $Cl_{K}$. Unfortunately, the action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in G$ and $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{\alpha}])$ will not be well-defined in general as $\sigma(\frac{1}{\alpha})=\sigma(\alpha)^{-1}$ need not be in $\mathcal{O}_{K}[\frac{1}{\alpha}]$ in general. However, in the case that $\alpha\in\mathbb{Z}$, we avoid this issue as we will have $\sigma(\frac{1}{\alpha})=\frac{1}{\alpha}$ for all $\sigma\in G$.
Corollary 4.4.
Let $K$ be a Galois number field and $A$ the subgroup of $\text{Aut}(Cl_{K})$ induced by the action of of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$. Suppose that $\alpha\in\mathcal{O}_{K}$ is a nonzero nonunit and
$(\alpha)=\mathfrak{p}_{1}^{n_{1}}\mathfrak{p}_{2}^{n_{2}}\dotsm\mathfrak{p}_{k}^{n_{k}}.$
Then $Cl(\mathcal{O}_{K}[\frac{1}{N(\alpha)}])\cong Cl_{K}/A\left<[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{k}]\right$> where $N(\alpha)$ is the norm of $\alpha$, and $A\left<\alpha\right$> denotes the subgroup generated by the $A$-automorphic images of the elements in $\alpha$.
Proof.
Recall that the ideal norm agrees with the element norm for principal ideals. Thus, as $K$ is Galois, we have $(N(\alpha))=N((\alpha))=\prod_{\sigma\in G}\sigma((\alpha))=\prod_{\sigma\in G}\sigma(\mathfrak{p}_{1})^{n_{1}}\dotsm\sigma(\mathfrak{p}_{k})^{n_{k}}$. The theorem then follows by applying Theorem 4.1 to $N(\alpha)$. ∎
Now, the group action of $G$ on the class group of $\mathcal{O}_{K}[\frac{1}{N(\alpha)}]$ will be well-defined. More than this, it will be a norm-like Galois action.
Theorem 4.5.
For a Galois number field $K$ and nonzero rational integer $n$, the group action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in\text{Gal}(F/\mathbb{Q}),[\mathfrak{P}]\in Cl\left(\mathcal{O}_{K}[\frac{1}{n}]\right)$ is a well-defined, norm-like action.
Proof.
It is sufficiently clear that conditions 1-3 from Definition 1.1 hold. Now, let $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{n}])$ where $\mathfrak{P}$ is a prime ideal representative. Note, such a $\mathfrak{P}$ must exist because each class of $Cl_{K}$ contains infinitely many primes and only finitely many intersect $\{\alpha^{n}\,|\,n\in\mathbb{N}$}. Thus, existence follows from the 1-1 correspondence between prime ideals of $\mathcal{O}_{K}[\frac{1}{\alpha}]$ and prime ideals of $\mathcal{O}_{K}$ disjoint from $S=\{\alpha^{n}\,|\,n\in\mathbb{N}\}$ given by $\mathfrak{p}\longleftrightarrow\mathfrak{p}\mathcal{O}_{K}[\frac{1}{\alpha}]$ for $\mathfrak{p}\in\text{Spec}(\mathcal{O}_{K})$.
If we write $\mathfrak{p}=\mathfrak{P}\cap\mathcal{O}_{K}$, we claim that $\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{n}]=\prod_{\sigma\in G}\sigma(\mathfrak{P})$, and thus the norm property (4) is also inherited from our original group action. We begin by noting that for any $\sigma\in G$, $\sigma(\mathcal{O}_{K})\subseteq\mathcal{O_{K}}$. Now, taking $\sigma^{-1}$ on both sides, we see that $\mathcal{O}_{K}\subseteq\sigma^{-1}(\mathcal{O}_{K})$ for any $\sigma\in G$, so in fact $\sigma(\mathcal{O}_{K})=\mathcal{O}_{K}$. Now, writing $G=\{\sigma_{i}\}_{i=1}^{n}$, let $x\in\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{n}]$ and write $x=\left(\prod_{i=1}^{n}\sigma_{i}(p_{i})\right)\cdot\frac{a}{n^{k}}$ where $p_{i}\in\mathfrak{p}$ and $a\in\mathcal{O}_{K}$. Let $b\in\mathcal{O}_{K}$ such that $\sigma(b)=a$. Then, $x=\sigma_{1}(\frac{p_{1}b}{n^{k}})\left(\prod_{i=2}^{n}\sigma_{i}(p_{i})\right)\in\prod_{\sigma\in G}\sigma(\mathfrak{P})$ as $\frac{p_{1}b}{n^{k}}\in\mathfrak{p}_{S}=\mathfrak{P}.$ For the other containment, let $y\in\prod_{\sigma\in G}\sigma(\mathfrak{P})$, and write $y=\prod_{i=1}^{n}\sigma_{i}(\frac{\bar{p}_{i}}{n^{r_{i}}})=\left(\prod_{i=1}^{n}\sigma_{i}(\bar{p}_{i})\right)\cdot\prod_{i=1}^{n}\frac{1}{n^{r_{i}}}\in\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{n}]$. ∎
The well-definedness of the action follows from the fact that $\sigma\left(\frac{1}{n}\right)=\frac{1}{n}\in\mathcal{O}_{K}[\frac{1}{n}]$ and thus $\sigma(\mathcal{O}_{K}[\frac{1}{n}])=\mathcal{O}_{K}[\frac{1}{n}]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$. More on this in the next section. Now, any restrictions the structure of $G$ places on $Cl_{K}$ must also hold for $Cl(\mathcal{O}_{K}[\frac{1}{n}])$. In particular, on $Cl(\mathcal{O}_{K}[\frac{1}{N(\alpha)}]$ for any nonzero, nonunit $\alpha$. Thus, by Corollary 4.4, if the set $\{[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\}$ is closed under the action of $G$, then we may apply our previous methods to $Cl_{K}/\left<[\mathfrak{p}_{1}],...,[\mathfrak{p}_{k}]\right$>. This is just another example of how misleading the quadratic case can be. As $\text{Gal}(K/\mathbb{Q})$ induces the subgroup $\{id,-id\}\leqslant\text{Aut}(Cl_{K})$ where $id$ is the identity and $id([I])=[I]^{-1}$, we see that any homomorphic image of $Cl_{K}$ can be realized by localizing at an integer—allowing significant restriction on $Cl_{K}$.
In general, it will be sufficient for a subgroup to be $\text{Aut}(Cl_{K})$-invariant. For example, if $G$ cannot admit a Galois action on $\mathbb{Z}/n\mathbb{Z}$, then we cannot have $Cl_{K}\cong\mathbb{Z}/an\mathbb{Z}$ for any $a\in\mathbb{N}$. This follows from the fact that $\mathbb{Z}/an\mathbb{Z}$ has a unique subgroup of order $n$. With this in hand, Theorem 2.4 becomes an immediate corollary of Theorem 3.1 in [12]. Furthermore, we can quickly improve upon Theorem 3.1.
Corollary 4.6.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, $Cl_{K}\not\cong\mathbb{Z}/np^{2}\mathbb{Z}$ for any $n\in\mathbb{N}$.
Once again, as we must have $h_{K}\equiv 0$ or $1$ mod $p$, this constitutes a serious restriction for extensions of (odd) prime degree. Now, note also that any Sylow $q$-subgroup of $Cl_{K}$ is invariant under automorphism. Thus, by the same logic, any restriction placed on $Cl_{K}$ may also be applied to its Sylow $q$-subgroups. Applying this also to Theorem 3.1, the following stronger result which appears in [21] is immediate.
Corollary 4.7.
Let $K$ be a Galois number field of degree $p$ where $p$ is an odd prime. Then, the Sylow $p$-subgroup of $Cl_{K}$ is not isomorphic to $\mathbb{Z}/p^{n}\mathbb{Z}$ for any $n\geqslant 2$.
In the same way, we can greatly strengthen Theorem 2.2.
Corollary 4.8.
Let $K$ be a Galois number field of degree $p^{r}$ and $S(q)$ a non-trivial Sylow $q$-subgroup of $Cl_{K}$. Then, $p=q$ or $|S(q)|\equiv 1$ mod $p$. Therefore, if $h_{k}=p^{r}q_{1}^{n_{1}}\dotsm q_{k}^{n_{k}}$, then $q_{i}^{n_{i}}\equiv 1$ mod $p$ for all $1\leqslant i\leqslant k$.
This was proven for abelian number fields by Fröhlich ([14]) which is a notably more restrictive condition. Now, observe that while Theorem 2.2 would not preclude an extension of degree $K:\mathbb{Q}]=3^{r}$ from having $h_{k}=55$, Corollary 4.8 shows this is impossible. In fact, such an extension cannot have $h_{K}=5n$ for any $n\in\mathbb{N}$ not divisible by $5$.
5. Overrings of Dedekind Domains
We conclude with a general consideration of overrings of Dedekind domains with eye toward developing a general criterion for when the class group of an overring of $\mathcal{O}_{K}$ is a Galois module via the action we have been studying. In particular, for an overring $\mathcal{O}_{K}\subseteq R\subseteq K$, we would like to know when the Galois group preserves fractional ideals of $R$ and thus admits the same norm-like group action on $Cl(R)$. We saw in Section 5 that even in the case of simple extensions, this may fail. Consider the following example.
Example 5.1.
Let $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-5}]$ and consider the overring $\mathcal{O}_{K}\left[\frac{1}{1+\sqrt{-5}}\right]$. Now, $\text{Gal}(K/\mathbb{Q})=\{\textbf{1},\sigma\}$ where $\sigma$ represents complex conjugation. Note that $\sigma(R)$ fails to be an $R$-module as $\frac{1}{1+\sqrt{-5}}\cdot\sigma\left(\frac{1}{1+\sqrt{-5}}\right)=\frac{1}{1+\sqrt{-5}}\cdot\frac{1}{1-\sqrt{-5}}=\frac{1}{6}\notin R$.
The following is a well-known characterization of overrings of Dedekind domains which first appeared as Theorem 1 in [16].
Proposition 5.2.
Let $D$ be a Dedekind domain and $R$ an overring of $D$. Then
$R=\underset{\mathfrak{p}R\neq R}{\bigcap}D_{\mathfrak{p}}$
where the $\mathfrak{p}$ range over $\text{Spec}(D)$.
The following result can be found in [17], and it follows quickly from the preceding proposition. We include a proof here as it is constructive and will inform later results and examples.
Theorem 5.3.
Let $D$ be a Dedekind domain with torsion class group. There exists a 111-1 correspondence between overrings of $D$ and nonempty subsets of $\text{Spec}(D)$ given by
$S\longleftrightarrow\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}.$
Proof.
By Proposition 5.2, every overring of $D$ can be expressed in the form $\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}$ for some $S\subseteq\text{Spec}(D)$. Thus, it suffices to show that $\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}\neq\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}$ for subsets $S_{1},S_{2}\subseteq\text{Spec}(D)$ whenever $S_{1}\neq S_{2}$. Without loss of generality, assume $\mathfrak{q}\in S_{1}\backslash S_{2}.$ Recall that $Cl(D)$ is torsion, so $|[\mathfrak{q}]|:=n$ is finite. Thus, $\mathfrak{q}^{n}=(q)$ for some $q\in D$. Now, $\mathfrak{q}$ is the unique prime ideal containing $q$, so $\frac{1}{q}\in\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}\backslash\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}$. ∎
Corollary 5.4.
Let $D$ be a Dedekind domain with torsion class group. Then $D\subseteq D[\frac{1}{a}]$ is a minimal ring extension if and only if $aD=\mathfrak{p}^{k}$ for some $k\in\mathbb{N}$.
Proof.
We prove the forward implication by contraposition. Assume there are distinct primes $\mathfrak{p},\mathfrak{q}\in\text{Spec}(D)$ which divide $aD$. Then, $a\in\mathfrak{p},\mathfrak{q}$, and hence $\mathfrak{p}D[\frac{1}{a}]=\mathfrak{q}D[\frac{1}{a}]=D$. Thus, by Corollary 5.3, there is an overring properly between $D$ and $D[\frac{1}{a}]$.
Now, for the reverse implication, assume $aD=\mathfrak{p}^{k}$ and $\mathfrak{q}D[\frac{1}{a}]=D[\frac{1}{a}]$ for some prime $\mathfrak{q}\in\operatorname{Spec}(D)$. Because $D=\mathfrak{p}^{k}, aa$ is strongly irreducible in $D$, and thus the saturation of $\{a^{n}\,|\,n\in\mathbb{N}\}$ in $D$ is simply $\{a^{n}\,|\,n\in\mathbb{N}\}$. Hence, $a^{k}\in\mathfrak{q}$ for some $k\in\mathbb{N}$ which implies $a\in\mathfrak{q}$. However, $\mathfrak{p}$ is the unique prime ideal containing $a$, so $\mathfrak{p}=\mathfrak{q}$. Therefore, $\mathfrak{p}$ is the unique prime ideal of $D$ such that $\mathfrak{p}D[\frac{1}{a}]=D[\frac{1}{a}]$. By the correspondence in Corollary 5.3, we have the $D\subseteq D[\frac{1}{a}]$ must be a minimal extension.
By the correspondence in Theorem 5.3, we see that for any adjacent ring $R$ over D, there is a unique $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}R=R$. What is more, recalling that any adjacent extension simple, $\mathfrak{p}$ uniquely determines the irreducible $a\in D$ (up to a unit) such that $R=D[\frac{1}{a}]$. In particular, we have $(a)=\mathfrak{p}^{n}$ where $n=|[\mathfrak{p}]|$. Inspired by this uniqueness, we present the following Jordan–Hölder like result on sequences of adjacent overring.
Theorem 5.5.
Let $D$ be a Dedekind domain with finite group $Cl(D)$ such that every class contains a prime ideal. Let $R$ be an overring of $D$ which is minimal with respect to being a PID. If $D:=D_{0}\subseteq D_{1}\subseteq D_{2}\subseteq\dotsm\subseteq D_{r_{1}}:=R$ and $D:=D_{0}^{\prime}\subseteq D_{1}^{\prime}\subseteq D_{2}^{\prime}\subseteq\dotsm\subseteq D_{r_{2}}^{\prime}:=R$ are sequences of adjacent domains, then $r_{1}=r_{2}$ and the sequence of $\alpha_{i}\in R$ such that $D_{i+1}=D_{i}[\alpha_{i}]$ is unique up to reordering. What is more, such an overring is guaranteed to exists.
Proof.
We start by noting that for any adjacent domain $D\subseteq D_{1}$ there exists a unique $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}D_{1}=D_{1}$ by Theorem 5.3. Thus, we must have $D_{1}=\cap_{\mathfrak{q}\neq\mathfrak{p}}D_{\mathfrak{q}}=D[\frac{1}{a}]$ where $aD=\mathfrak{q}^{n}$ and $n=|[\mathfrak{q}]|$. Recall that $D[\frac{1}{a}]$ is the localization of $D$ at the complement of the set-theoretic union of all prime ideals of $D$ not containing $a$. In this case, that is all prime ideals omit $\mathfrak{q}$. By continuing this process and recalling the correspondence between prime ideals of a localization $D_{S}$ and prime ideals of $D$ disjoint from $S$ given by $\mathfrak{p}\mapsto\mathfrak{p}D_{S}$, we see that there is a unique, finite set of $r_{1}$ primes $\mathfrak{p}_{1},...,\mathfrak{p}_{r_{1}}\in\text{Spec}(D)$ such that $\mathfrak{p}R=R$. Thus, any sequence of adjacent domains between $D$ and $R$ must have length $r_{1}:=r$ and is unique up to our ordering of the $\mathfrak{p}_{i}$. In particular, if $\mathfrak{p}_{i}D_{i-1}$ is the unique prime such that $\mathfrak{p}_{i}D_{i}=D_{i}$, we must have $D_{i}=D_{i-1}[\frac{1}{a_{i}}]$ where $a_{i}D_{i-1}=\mathfrak{p}_{i}^{n_{i}}$.
Now, to construct such a PID and sequence, take some non-principal $\mathfrak{q}_{1}\in D$ such that $\mathfrak{q}_{1}^{n_{1}}=b_{1}D$, and let $D_{1}=D[\frac{1}{b}]$. As $\mathfrak{q}_{1}]\neq\text{Prin}(D)$, we must have $|Cl(D_{1})|<|Cl(D)|$. By the finiteness of the class group, if we continue this process, we will eventually have $|Cl(D_{t})|=1$ for some $t\in\mathbb{N}$ with $|Cl(D_{i})|\neq 1$ for all $0<i<t$. Thus, $D_{t}$ will be minimal with respect ot being a PID. ∎
We will now shift our focus back to the particular case of rings of integers. For a more general treatment of overrings of Dedekind domains and their class groups, the reader is encouraged to see [9] and [17]. Theorem 5.5 tells us that any ring of integers in particular is finitely many adjacent steps away from being a PID. In this setting, this is equivalent to asking how far the ring is from being a UFD—a question considered more broadly in [2].
Example 5.6.
Recall $Cl_{K}\cong\mathbb{Z}/4\mathbb{Z}$ for $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-14}]$. Via the process described in the proof of Theorem 5.5, $\mathbb{Z}[\sqrt{-14}]\subseteq\mathbb{Z}[\sqrt{-14},\frac{1}{2}]\subseteq\mathbb{Z}[\sqrt{-14},\frac{1}{6}]$ is a series of adjacent domains with $\mathbb{Z}[\sqrt{-14},\frac{1}{6}]$ a minimal PID over $\mathbb{Z}[\sqrt{-14}]$.
Now, even in the case of rings of integers, $R$ being a minimal PID over $\mathcal{O}_{K}$ does not imply the existence of such chain of adjacent domains. In fact, the following example implies that there always exists a PID over $\mathcal{O}_{K}$ failing to have this property except in the case that the ring of integers itself is a PID.
Example 5.7.
Let $\mathcal{O}_{K}$ be a ring of integers with $Cl_{K}=\mathbb{Z}/2\mathbb{Z}$, and let $P$ be the set of all non-principal prime ideals of $\mathcal{O}_{K}$. Now, taking $S=\cup_{\mathfrak{p}\in P}\mathfrak{p}$, we have that $R_{S}=\{\frac{r}{a}\,|\,r\in R,a\notin S\}$ also has class group $\mathbb{Z}/2\mathbb{Z}$ by Theorem 1.5 of [7]. Note, this is equivalent to localizing at the multiplicatively closed set generated by the prime elements of $\mathcal{O}_{K}$. Now, take a (non-principal) prime $\mathfrak{q}\in\text{Spec}(R)$, and let $aR_{S}=\mathfrak{q}^{n}$ for some nn\ $\in\mathbb{N}$. Then, $R_{S}[\frac{1}{a}]$ is minimal over $\mathcal{O}_{K}$ with respect to being a PID, but there is no finite chain of adjacent domains between them by Theorem 5.3.
We now return to the particular case of rings of integers. For a more general treatment of overrings of Dedekind domains, the reader is referred to [7] and [8]. In the following theorem, we adopt the notation of [24] and denote the group of fractional ideals of a Dedekind domain $R$ as $G(R)$.
Theorem 5.8.
Let $K$ be a Galois number field with $G:=\text{Gal}(K/\mathbb{Q})$ and $R$ an overring of the ring of integers $\mathcal{O}_{K}$. The following are equivalent:
  1. (1)
    The action $\sigma\cdot[I]=[\sigma(I)]$ for $\sigma\in G$ and $I]\in G(R)$ is well-defined.
  2. (2)
    $R$ is Galois-invariant. That is, for all $\sigma\in G$, $\sigma(R)=R$.
  3. (3)
    For all $\mathfrak{p}\in\text{Spec}(\mathcal{O}_{K})$ and $\sigma\in G$, $\mathfrak{p}R=R\implies\sigma(\mathfrak{p})R=R$.
  4. (4)
    For any $\mathfrak{p},\mathfrak{q}\in\text{Spec}(\mathcal{O}_{K})$ lying over the same rational prime $p$, $\mathfrak{p}R=R\implies\mathfrak{q}R=R$.
Proof.
$(1\Rightarrow 2)$ By well-definedness, we must have $\sigma(R)$ an $R$-module for all $\sigma\in G$. Then, for all $r\in R$, $r\cdot\sigma(1)=r\cdot 1=r\in\sigma(R)$, so we have $R\subseteq\sigma(R)$. By the same argument, $R\subseteq\sigma^{-1}(R)$ and thus $\sigma(R)\subseteq R$. Therefore, $\sigma(G)=G$ as desired.
$(2\Rightarrow 1)$ Assume $R$ is Galois-invariant. Let $I\in G(R)$ and $\sigma\in G$. By definition, we have $\alpha I\subseteq R$ for some $\alpha\in K$. By our assumption, $\sigma(\alpha)\sigma(I)=\sigma(\alpha I)\subseteq\sigma(R)\subseteq R$ for $\sigma(\alpha)\in K$. Now, let $r\in R$, then $r^{\prime}:=\sigma^{-1}(r)\in R$ by assumption. Therefore, $r\sigma(I)=\sigma(r^{\prime})\sigma(I)\in\sigma(r^{\prime}I)\subseteq\sigma(I)$ as $I$ is an $R$-module. We conclude that $\sigma(I)$ is a fractional ideal, and thus the map $I\mapsto\sigma(I)$ from $G(\mathcal{O}_{K})$ to $G(R)$ descends to the group action in $1$.
$(2\Rightarrow 3)$ We will proceed by contraposition. Assume that there exists $\mathfrak{p}\in\text{Spec}(\mathcal{O})_{K}$ and $\sigma\in G$ such that $\mathfrak{p}R=R$ but $\sigma(\mathfrak{p})R\neq R$. Then there exists $\frac{a}{b}\in U(R)\cap\mathfrak{p}$, so $\frac{b}{a}\in R$. However, $\sigma(\frac{b}{a})=\frac{\sigma(b)}{\sigma(a)}\notin R$ because $\sigma(\frac{a}{b})=\frac{\sigma(a)}{\sigma(b)}\in\mathfrak{q}$ and $\mathfrak{q}\cap U(R)=\emptyset$. Therefore, $R$ is not Galois-invariant.
$(3\Rightarrow 2)$ We will once again proceed by contraposition. Assume there exists $\frac{a}{b}\in R$ with $a,b\in\mathcal{O}_{K}$ and $\sigma\in G$ such that $\sigma(\frac{a}{b})=\frac{\sigma(a)}{\sigma(b)}\notin R$. Recall that $R=\cap_{\mathfrak{p}R\neq R}(\mathcal{O}_{K})_{\mathfrak{p}}$. Then, there exists some $\mathfrak{q}\in\text{Spec}(R)$ such that $\mathfrak{q}R\neq R$ with $\sigma(b)\in\mathfrak{q}$. Now, $b=\sigma^{-1}(\sigma(b))\in\sigma^{-1}(\mathfrak{q}):=\mathfrak{q}^{\prime}$ implies $\mathfrak{q}^{\prime}R=R$. Thus, we have $\mathfrak{q}^{\prime}\in\text{Spec}(R)$ such that $\mathfrak{q}^{\prime}R=R$, but $\sigma(\mathfrak{q}^{\prime})R=\mathfrak{q}R\neq R$.
$(3\Leftrightarrow 4)$ The equivalence of these conditions follows directly from the transitive action of $G$ on the primes lying over a given rational prime (for details, see [22]). This completes the proof. ∎
This theorem demonstrates that the fact that any class group of an overring of $\mathcal{O}_{K}$ which admits such a group action can be realized an a simple extension of the form $\mathcal{O}_{K}[\frac{1}{n}]$ where $n$ is an integer. In particular, we see that the types of extensions studied in the previous section are sufficient to achieve all homomorphic images which are of use to us in restricting the structure of $Cl_{K}$. The theorem also provides a general framework for using this approach to understanding particular class groups. Consider the following corollary and example.
Corollary 5.9.
Let $K$ be a Galois number field of degree $p^{n}$ and $q$ a rational prime. If $H$ is the subgroup of $Cl_{K}$ generated by the classes of primes lying over $q$, then $Cl_{K}:H]\equiv 0$ or $1$ mod $p$.
Proof.
Theorem 4.5 tells us that $\text{Gal}(K/\mathbb{Q})$ admits a norm-like action on the class group of $\mathcal{O}_{K}[\frac{1}{q}]$ which will have order $Cl_{K}:H]$. The result then follows directly from Theorem 2.2. ∎
In the case of Galois number fields of prime degree, Corollary 5.9 tells us that we can glean much about the class group by considering the cyclic subgroups generated by the unique prime ideals lying over ramified primes.
6. The Arithmetic of the Normset
In this section, we consider some arithmetic implications of the action we have been considering. We will begin by developing the notion of the Davenport constant of a weighted zero-sum sequence monoid from a number theoretic point of view, leading to a nice arithmetic interpretation as well as some illustrative and novel examples. We conclude by making a connection between the arithmetic of quadratic norms and the partition problem—a classic NP-complete problem.
If $K$ is a Galois number field, the action we have been studying plays a central role in understanding the arithmetic of the normset $N_{K}:=\{|N(\alpha)|\,:\,\alpha\in\mathcal{O}_{K}^{*}\}$. This connection was first made in [11] where the author proved that $N_{K}$ is saturated if and only if $\text{{Gal}}(K/\mathbb{Q})$ acts trivially on $Cl_{K}$. This connection was greatly strengthened with the advent of weighted Davenport constants. Let $G$ be a finite abelian group. Recall that the Davenport constant of $G$, denoted $D(G)$, is the minimum natural number $n$ such that any $G$-sequence $\{g_{1},g_{2},...,g_{n}\}$ of (not necessarily distinct) elements of $G$ must have a zero-sum subsequence—that is, a subsequence summing to the identity. Equivalently, we may define it as the largest natural number $m$ such that there exists a zero-sum sequence $\{g_{1},g_{2},...,g_{m}\}$ with no proper zero-sum subsequence.
Now, for a finite abelian group $G$ and $\Gamma$ a set which operates on $G$ by endomorphism, $D_{\Gamma}(G)$ is the smallest positive integer $n$ such that given any $G$-sequence $\{g_{1},g_{2},...,g_{n}\}$, there exists a subsequence $\{g_{i_{1}},g_{i_{2}},...,g_{i_{k}}\}$ and weights $\sigma_{i_{1}},\sigma_{i_{2}},...,\sigma_{i_{n}}\in\Gamma$ such that $\{\sigma_{i_{1}}(g_{i_{2}}),\sigma_{i_{2}}(g_{i_{2}}),...,\sigma_{i_{k}}(g_{i_{k}})\}$ is a zero-sum sequence. Note that “operating by endomorphism” means precisely meeting conditions 3 of Definition 1.1 and thus inducing a map into $\text{End}(G)$. We refer to $D_{\Gamma}$ as a weighted Davenport constant. This definition closely mimics the the definition of $D(G)$, and in fact $D_{\{1\}}(G)=D(G)$. What is more, just as the Davenport constant is intimately related to the arithmetic of $\mathcal{O}_{K}$, the weighted Davenport constant is closely related to that of $N_{K}$.
Weighted Davenport constants were first studied by Adhikari et al in [1], but it was Halter-Koch who first made the connection with the ideal arithmetic of rings of integers in [18]. In particular, if we allow $\Gamma=\text{Gal}(K/\mathbb{Q})$ for a Galois number field $K$, Theorem 2.2 of [18] gives a nice characterization of $D_{\Gamma}(Cl_{K})$ in terms of norms of ideals in $\mathcal{O}_{K}$.
There has been a recent resurgence of interest in this area through the study of monoids of weighted zero-sum sequences ([5], [15]). Given a subset $\Gamma\subseteq\text{End}(G)$, a $G$-sequence $\{g_{1},g_{2},...,g_{n}\}$ is called a $\Gamma$-weighted zero-sum sequence if there exist $\sigma_{1},\sigma_{2},...,\sigma_{n}\in\Gamma$ such that $\{\sigma_{1}(g_{1}),\sigma_{2}(g_{2}),...,\sigma_{n}(g_{n})\}$ is a zero-sum sequence. The set of all weighted zero-sum sequences forms a monoid under concatenation which is denoted $\mathfrak{B}_{\Gamma}(G)$. With this notation, we see that $D_{\Gamma}(G)$ is the minimal $n$ such that any $G$-sequence must have a $\Gamma$-weighted zero-sum subsequence.
A $\Gamma$-weighted zero-sum sequence $\{g_{1},g_{2},...,g_{n}\}$ is called minimal if it cannot be partitioned into two nonempty $\Gamma$-weighted zero-sum sequences. Then $D(\mathfrak{B}_{\Gamma}(G))$ denotes the largest positive integer $m$ such that there exists a minimal $\Gamma$-weighted zero-sum sequence of length $m$. Note that a $\Gamma$-weighted zero-sum sequence $\{g_{1},g_{2},...,g_{n}\}$ may be minimal and have a proper $\Gamma$-weighted zero-sum subsequence. Hence, we see that $D_{\Gamma}(G)$ and $D(\mathfrak{B}_{\Gamma}(G))$ are related but distinct.
For the remainder of this paper, we will focus on the case where $\Gamma:=\text{Gal}(K/\mathbb{Q})$ for a Galois number field $K$. In this case, the authors of [5] showed that there exists a transfer homomorphism from $N_{K}$ to $\mathfrak{B}_{\Gamma}(Cl_{K})$. In essence, this means that factorization properties of the normset are completely characterized by the arithmetic of these $\Gamma$-weighted zero-sum sequences—that is, by the action of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$. Extensive results on $D(\mathfrak{B}_{\Gamma}(Cl_{K}))$ and implications for the factorization properties of $N_{K}$ are studied through this lens in [5] and [15]. We will presently develop the connection between these two objects from a strictly number-theoretic point of view. This will allow us to give a nice arithmetic interpretation of $D(\mathfrak{B}_{\Gamma}(Cl_{K}))$ and provide some enlightening new examples.
First, let us recall that for a Galois number field $K$ and $\alpha\in\mathcal{O}_{K}$, $N(\alpha)=\prod_{\sigma\in\text{Gal}(K/\mathbb{Q})}\sigma(\alpha)$. Thus, if we write $(\alpha)=\mathfrak{p}_{1}\mathfrak{p}_{2}\dotsm\mathfrak{p}_{m}$ where the prime ideals are not necessarily distinct, it is not hard to see that $N(\beta)=N(\alpha)$ if and only if there exist $\bar{\sigma}_{1},\bar{\sigma}_{2},...,\bar{\sigma}_{m}\in\text{Gal}(K/\mathbb{Q})$ such that $(\beta)=\bar{\sigma}_{1}(\mathfrak{p}_{1})\bar{\sigma}_{2}(\mathfrak{p}_{2})\dotsm\bar{\sigma}_{m}(\mathfrak{p}_{m})$. This follows from the uniqueness of prime ideal factorization as $N(\alpha)$ agrees with the ideal norm $N((\alpha))=\prod_{\sigma\in\text{Gal}(K/\mathbb{Q})}\sigma((\alpha))=\prod_{\sigma\in\text{Gal}(K/\mathbb{Q})}\sigma(\mathfrak{p}_{1})\sigma(\mathfrak{p}_{2})\dotsm\sigma(\mathfrak{p}_{m})$. Note that $\{[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],...,[\mathfrak{p}_{m}]\}$ is a zero-sum sequence in $Cl_{K}$, so finding such an element $\beta$ corresponds to finding weights $\phi_{1},\phi_{2},...,\phi_{m}$ in the subgroup of $\text{Aut}(Cl_{K})$ induced by the action of $\text{Gal}(K/\mathbb{Q})$ such that $\{\phi_{1}[\mathfrak{p}_{1}],\phi_{1}[\mathfrak{p}_{1}],...,\phi_{n}[\mathfrak{p}_{m}]\}$ is also a zero-sum sequence. We will consider the problem of such norm collisions in more detail later. Let us write $\text{Gal}(K/\mathbb{Q})=\{\sigma_{i}\}_{i=1}^{n}$ and consider the table
[…]
$\vdots$ $\vdots$ $\vdots$ $\ddots$ $\vdots$
[…]
containing all the factors of the ideal norm $N((\alpha))$. Once again, we can determine all other elements with norm $N(\alpha)$ (up to a unit multiple) by selecting one prime from each row such that the resulting product is principal. Now, observe that $N(\alpha)$ will be reducible in $N_{K}$ if and only if there exist $\sigma_{j_{k}}\in\text{Gal}(K/\mathbb{Q})$ such that $\sigma_{j_{1}}(\mathfrak{p}_{1})\sigma_{j_{2}}(\mathfrak{p}_{2})\dotsm\sigma_{j_{n}}(\mathfrak{p}_{m})=(\beta)$ where $\beta$ is reducible in $\mathcal{O}_{K}$. That is, $\beta=\gamma_{1}\gamma_{2}$ where $\gamma_{1},\gamma_{2}$ are nonunits, so $(\beta)=(\gamma_{1})(\gamma_{2})=\sigma_{j_{1}}(\mathfrak{p}_{1})\sigma_{j_{2}}(\mathfrak{p}_{2})\dotsm\sigma_{j_{n}}(\mathfrak{p}_{m})$. By uniqueness of prime factorization, this corresponds to partitioning $\{\mathfrak{p}_{1},\mathfrak{p}_{2},...,\mathfrak{p}_{m}\}$ into two nonempty $\Gamma$-weighted zero-sum sequences. From this, we get the following characterization of $D(\mathfrak{B}_{\Gamma}(Cl_{K}))$.
Proposition 6.1.
Let $K$ be a Galois number field and $\Gamma=\text{Gal}(K/\mathbb{Q})$. If $\alpha\in\mathcal{O}_{K}$ with $N(\alpha)$ irreducible in $N_{K}$, then $D(\mathfrak{B}_{\Gamma}(Cl_{K}))$ is a sharp upper bound for the length of the prime factorization (counting multiplicity) of $(\alpha)$.
This is especially nice as it mirrors the classic characterization of $D(Cl_{K})$ as a sharp upper bound for the length of the prime factorization of $(\alpha)$ for any irreducible element $\alpha\in\mathcal{O}_{K}$.
Example 6.2.
Consider $K=\mathbb{Q}(\sqrt{-14})$ and $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-14]}$ with $Cl_{K}\cong\mathbb{Z}/4\mathbb{Z}$ generated by $\mathfrak{p}]$ where $\mathfrak{p}=(3,1-\sqrt{-14})$. Now, $\{[\mathfrak{p}],[\mathfrak{p}],[\mathfrak{p}],[\mathfrak{p}]\}$ is a minimal zero-sum sequence of $Cl_{K}$, so $\mathfrak{p}^{4}=(5-2\sqrt{-14})$ implies $5-2\sqrt{-14}$ is irreducible in $\mathbb{Z}[\sqrt{-14}]$. Now, $G=\text{Gal}(K/\mathbb{Q})=\{\sigma_{1},\sigma_{2}\}$ where $\sigma_{1}$ is the identity and $\sigma_{2}$ is complex conjugation. Thus, $\sigma(\mathfrak{p})=\mathfrak{p}$ and $\sigma_{2}(\mathfrak{p})=\bar{\mathfrak{p}}=(3,1+\sqrt{-14})$. Now, let us make our selections.
$\boxed{\sigma_{1}(\mathfrak{p})}$ $\sigma_{1}(\mathfrak{p})$ $\boxed{\sigma_{1}(\mathfrak{p})}$ $\sigma_{1}(\mathfrak{p})$
$\sigma_{2}(\mathfrak{p})$ $\boxed{\sigma_{2}(\mathfrak{p})}$ $\sigma_{2}(\mathfrak{p})$ $\boxed{\sigma_{2}(\mathfrak{p})}$
Now, $\sigma_{1}(\mathfrak{p})\sigma_{2}(\mathfrak{p})\sigma_{1}(\mathfrak{p})\sigma_{2}(\mathfrak{p})=(\mathfrak{p}\bar{\mathfrak{p}})^{2}=(3)^{2}=(9)$. Therefore, we must have $N(5-2\sqrt{-14})=N(9)=N(3)^{2}$, and we see in fact $N(5-2\sqrt{-14})=(5-2\sqrt{-14})(5+2\sqrt{-14})=81=9^{2}=N(3)^{2}$. So we see that while $5-2\sqrt{-14}$ was irreducible, $N(5-2\sqrt{-14})$ need not be.
A significant majority of the work on $D_{\Gamma}(G)$ and $D(\mathfrak{B}_{\Gamma}(G))$ has been in the plus-minus case. That is, where the set of weights is taken to be $\{id,-id\}$ where id(g)=g-id(g)=-g. For our context, this corresponds to the case where $K$ is a quadratic ring of integers as $\text{Gal}(K/\mathbb{Q})=\{id,\sigma\}$ where $\sigma$ represents complex conjugation and acts on $Cl_{K}$ by sending any element to its inverse. This follows directly from the transitive action of the Galois group on primes lying over rational primes. We will now consider a particularly nice example outside of this family.
Example 6.3.
Let $K=\mathbb{Q}(\zeta_{37})$. Famously, we have $Cl_{K}\cong\mathbb{Z}/37\mathbb{Z}$. Now consider the rational prime $149$. As $149\equiv 1$ mod $37$, we must have $(149)$ split into $36$ distinct primes in $K$. Remarkably, it can be shown that each of these distinct primes also lies in a distinct class of $Cl_{K}$. Hence, $\text{Gal}(K/\mathbb{Q})$ must act transitively on all the non-trivial elements of $Cl_{K}$. This means for any minimal zero-sum sequence $\{[\mathfrak{p}_{1}],[\mathfrak{p}_{2}],[\mathfrak{p}_{3}],[\mathfrak{p}_{4}]\}$ in $Cl_{K}$, there exist $\sigma_{2},\sigma_{4}\in\text{Gal}(K/\mathbb{Q})$ such that $\{[\mathfrak{p}_{1}],\sigma_{2}\cdot[\mathfrak{p}_{2}],[\mathfrak{p}_{3}],\sigma_{4}\cdot[\mathfrak{p}_{4}]\}=\{[\mathfrak{p}_{1}],[\mathfrak{p}_{2}]^{-1},[\mathfrak{p}_{3}],[\mathfrak{p}_{4}]^{-1}\}$ (note that we cannot have any $\mathfrak{p}_{i}]$ trivial because the original zero-sum sequence was minimal). This implies $D(\mathfrak{B}_{\Gamma}(Cl_{K}))\leqslant 3$. The sequence $\{[\mathfrak{p}^{12}],[\mathfrak{p}^{12}],[\mathfrak{p}^{12}]\}$ where $\mathfrak{p}$ is any non-principal prime shows us that we actually have equality: $D(\mathfrak{B}_{\Gamma}(Cl_{K}))=3$. Note that in this case $D_{\Gamma}(Cl_{K})=2\neq D(\mathfrak{B}_{\Gamma}(Cl_{K}))$. Now, Theorems 5.7 and 7.1 in [5] tell us that $\rho(N_{K})=D(\mathfrak{B}_{\Gamma}(Cl_{K}))/2=\frac{3}{2}$ while it is well-known that $\rho(\mathcal{O}_{K})=\frac{D(Cl_{K})}{2}=\frac{37}{2}$.
This is a shocking disparity, especially considering $N_{K}$ is half-factorial if and only if $\mathcal{O}_{K}$ is half-factorial for a Galois number field $K$ ([13]). Heuristically, we would expect the disparity between these two values to increase as the size of the induced subgroup of $\text{Aut}(Cl_{K})$ does. In the case of a quadratic number field with class group $\mathbb{Z}/2n\mathbb{Z}$ and $n\geqslant 2$, we have $D(\mathfrak{B}_{\Gamma}(Cl_{K}))=1+\frac{n}{2}$ ([5]), so $\frac{\rho(\mathcal{O_{K}})}{\rho(N_{K})}=\frac{2n}{n+2}$ and the order of the induced subgroup is $2$. For $K=\mathbb{Q}(\zeta_{37})$, we have $\frac{\rho(\mathcal{O_{K}})}{\rho(N_{K})}=\frac{37}{3}$ and the order of the induced subgroup is $36$. These observations motivate the following conjecture.
Conjecture 6.4.
Let $p_{i}$ be the $i^{th}$ irregular prime and $K=\mathbb{Q}(\zeta_{p_{i}})$
$\lim_{i\to\infty}\frac{\rho(\mathcal{O}_{K})}{\rho(N_{K})}=\infty.$
Connections have recently been made between the arithmetic of this action and coding theory in [4]. We conclude by returning to quadratic number fields and drawing an interesting connection between the arithmetic of their normsets and a classic problem from computational complexity theory. In particular, we are interested in the non-trivial norm collisions mentioned earlier in this section. Let us formalize this concept as a decision problem which we will refer to as the equal norm problem.
Equal Norm Problem.
Let $K$ be a Galois number Field and a nonzero nonunit $\alpha\in\mathcal{O}_{K}$. The equal norm problem is to decide if there exists $\beta\in\mathcal{O}_{K}$ with $N(\alpha)=N(\beta)$ and $(\alpha)\neq(\beta)$.
Note that the condition $(\alpha)\neq(\beta)$ means $\beta$ is not a unit multiple of $\alpha$—precluding trivial solutions. We now recall a classic NP-complete decision problem.
Partition Problem.
Given a nonempty multiset $S$ of positive integers, the partition problem is to decide whether or not $S$ can be partitioned into two subsets $S_{1}$ and $S_{2}$ such that the sums of the numbers in $S_{1}$ and $S_{2}$ are equal.
Our goal now is to show that any instance of the partition problem can be realized as an instance of the equal norm problem. With this goal in mind, we present the following theorem which follows directly as a porism of Theorem 1 in [3].
Theorem 6.5.
For any positive integer $n$, there exists a square free integer $d$ such that $(3)=\mathfrak{p}\mathfrak{q}$ in $\mathcal{O}_{K}$ where $K=\mathbb{Q}(\sqrt{-d})$, and $\mathfrak{p}]$ has order $n$ in $Cl_{K}$.
For the details of how to find such a $d$, see [3]. This theorem is precisely what we need to make the reduction in question.
Theorem 6.6.
Any instance of the partition problem can be realized as an instance of the equal norm problem for some $\alpha$ in a quadratic number field $K$. What is more, we have a deterministic algorithm for finding $\alpha$ and $K$.
Proof.
Let $S=\{a_{1},a_{2},...,a_{k}\}$ be a nonempty multiset of positive integers and write $n=\sum_{i-1}^{k}a_{i}$. By Theorem 6.6, there exists a square free integer $d$ such that $(3)=\mathfrak{p}\mathfrak{q}$ in $\mathcal{O}_{K}$ where $K=\mathbb{Q}(\sqrt{-d})$, and $\mathfrak{p}]$ has order $n$ in $Cl_{K}$. Now, consider the zero-sum sequence $\{[\mathfrak{p}^{a_{1}}],[\mathfrak{p}^{a_{2}}],...,[\mathfrak{p}^{a_{k}}]\}$ and let $\mathfrak{q}_{i}$ be a prime representative of $\mathfrak{p}^{a_{i}}]$ for all $1\leqslant i\leqslant k$. Then, $\{[\mathfrak{q}_{1}],[\mathfrak{q}_{2}],...,[\mathfrak{q}_{k}]\}$ is also a zero-sum sequence.
Now, let $(\alpha)=\mathfrak{q}_{1}\mathfrak{q}_{2}\dotsm\mathfrak{q}_{k}$. All other elements (up to unit multiplication) with norm $N(\alpha)$ can be achieved by discerning which $\mathfrak{q}_{i}$ we can change to $\sigma(\mathfrak{q}_{i})$ and achieve a principal ideal, where $\sigma$ is complex conjugation. As $\sigma$ acts on $Cl_{K}$ by sending elements to their inverses, finding an element $\beta\in\mathcal{O}_{K}$ with $N(\alpha)=N(\beta)$ and $(\alpha)\neq(\beta)$ corresponds to changing a nonzero number of the $\mathfrak{p}_{i}]$ to $\mathfrak{p}_{i}]^{-1}$ such that, without loss of generality, $\{[\mathfrak{p}^{a_{1}}]^{-1},...,[\mathfrak{p}^{a_{t}}]^{-1},[\mathfrak{p}^{a_{t}}],...,[\mathfrak{p}^{a_{k}}]\}$ is also a zero-sum sequence. Assume for the purpose of contradiction that $(\alpha)=(\beta)$. Then, by uniqueness of prime ideal factorization, we have some $\sigma(\mathfrak{q}_{i})=\mathfrak{q}_{j}\implies[\mathfrak{q}_{i}]^{-1}=[\mathfrak{q}_{j}]\implies[\mathfrak{p}^{a_{i}}]^{-1}=[p^{a_{j}}]\implies[\mathfrak{p}^{a_{i}+a_{j}}]=\text{Prin}(\mathcal{O}_{K})=[\mathfrak{p}^{n}]$. Thus, we must have $a_{i}+a_{j}=n$, but this contradicts our assumption that $k\geqslant 3$. Thus, the decision for the equal norm problem for $\alpha\in\mathcal{O}_{K}$ is yes if and only if (without loss of generality) we can add some nonzero number of weights weights such that $\{[\mathfrak{p}^{a_{1}}]^{-1},...,[\mathfrak{p}^{a_{t}}]^{-1},[\mathfrak{p}^{a_{t}}],...,[\mathfrak{p}^{a_{k}}]\}$ is a zero-sum sequence. This will hold if and only if $a_{1}-\dotsm-a_{t}+a_{t+1}+\dotsm a_{k}\equiv 0$ mod $n$ $\iff-a_{1}-\dotsm-a_{t}+a_{t+1}+\dotsm+a_{k}=0\iff a_{t+1}+\dotsm+a_{k}=a_{1}+\dotsm+a_{t}$ which we observe holds (without loss of generality) if and only if the decision for the partition problem for $S$ is yes. Finally, the case when $k\leqslant 1$ is degenerate, so we only need concern ourselves with $k=2$. As our only options are yes if $S=\{a,a\}$ and no if $S=\{a,b\}$ where $a=b$, we can pick an arbitrary affirmative and negative instance of the equal norm problem to assign in either case. This completes the proof. ∎
It is important to note that this reduction likely cannot be made in polynomial time because finding prime ideals representatives of the $\mathfrak{p}^{a_{i}}]$ classes is very likely a hard problem. Regardless, this theorem creates an interesting bridge between computational complexity and the arithmetic of Galois actions.
Example 6.7.
Consider the multiset $S=\{1,1,2,3,5\}$ with $1+1+2+3+5=12$. By [3], we want to find a square-free integer $d$ of the form $d=3^{12}-x^{2}$ for some integer $x$ with $2|x$ and $0<x<(2\cdot 3^{11})^{\frac{1}{2}}$, and we observe that $d:=3^{12}-2^{2}=531437=17\cdot 43\cdot 747$ suffices. Now, let $K=\mathbb{Q}(\sqrt{-d})$ so that $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-d}]:=\mathbb{Z}[\omega]$.
Now, $(3)=\mathfrak{p}_{1}\mathfrak{p}_{2}=(3,1+\omega)(3,2+\omega)$. We will take $\mathfrak{p}_{1}=(3,1+\omega)$, and we would like to find prime representatives of the classes $\{[\mathfrak{p}_{1}^{1}],[\mathfrak{p}_{1}^{1}],[\mathfrak{p}_{1}^{2}],[\mathfrak{p}_{1}^{3}],[\mathfrak{p}_{1}^{5}]\}$. We choose $\mathfrak{q}_{1}=\mathfrak{p}_{1}$, $\mathfrak{q}_{2}=(59077,59061+\omega)$, $\mathfrak{q}_{3}=(21379,21165+\omega)$, and $\mathfrak{q}_{5}=(3167,488+\omega)$ so that $\{[\mathfrak{p}_{1}^{1}],[\mathfrak{p}_{1}^{1}],[\mathfrak{p}_{1}^{2}],[\mathfrak{p}_{1}^{3}],[\mathfrak{p}_{1}^{5}]\}=\{[\mathfrak{q}_{1}],[\mathfrak{q}_{1}],[\mathfrak{q}_{2}],[\mathfrak{q}_{3}],[\mathfrak{q}_{5}]\}$ with each $\mathfrak{q}_{i}$ prime. We now compute $\mathfrak{q}_{1}\mathfrak{q}_{1}\mathfrak{q}_{2}\mathfrak{q}_{3}\mathfrak{q}_{5}=(35999493737049,23529598676827+\omega):=(\alpha)$.
Now, the equal norm problem for $\alpha$ will be equivalent to the partition problem for $S$. Observe that the decision for the partition problem is yes with solution $S_{1}=\{1,5\}$ and $S_{2}=\{1,2,3\}$. Therefore, $\bar{\mathfrak{p}}_{1}\mathfrak{p}_{1}\mathfrak{p}_{2}\mathfrak{p}_{3}\bar{\mathfrak{p}}_{5}=(\beta)$ for some $\beta\in\mathbb{Z}[\omega]$ where $\bar{\mathfrak{p}}=\{\bar{\gamma}\,|\,\gamma\in\mathfrak{p}\}$, and we have $N(\alpha)=N(\beta)$ with $(\alpha)\neq(\beta)$.